### v2

# Quiver moduli and quantum loop algebras

Abstract.
A classic result of Hernandez-Leclerc and Kashiwara-Kim-Oh-Park relates the $q$-characters of so-called reachable simple modules of quantum affine algebras to the Euler characteristics of certain quiver moduli spaces. We categorify and generalize this relation to all simple modules in the Hernandez-Jimbo category $\mathcal{O}$} using critical $K$-theory (our results hold for quantum toroidal as well as quantum affine algebras).
1. Introduction
1.1. The setting
Let ${\mathbb{K}}={\mathbb{Q}}(q)$ and fix a symmetrizable Kac-Moody Lie algebra $\mathfrak{g}$}, with a set $I$ of simple roots. We write $(d_{ij})_{i,j\in I}$ for the symmetrized generalized Cartan matrix of $\mathfrak{g}$}, see (63). The representation theory of the quantum loop algebra
$U_{q}(L{\mathfrak{g}})={\mathbb{K}}\Big\langle e_{i,d},f_{i,d},\varphi_{i,d^{\prime}}^{\pm}\Big\rangle_{i\in I,d\in{\mathbb{Z}},d^{\prime}\geq 0}\Big/\text{relations}$
has long been studied, see for instance [8, 25] for $\mathfrak{g}$} of finite type and [38] for $\mathfrak{g}$} of general type. We will consider the simple modules $L({\boldsymbol{\psi}})$ of the Borel subalgebra of $U_{q}(L{\mathfrak{g}})$ (defined in loc. cit.) which are indexed by so-called highest $\ell$-weights
(1) ${\boldsymbol{\psi}}=\left(\psi_{i}(z)=\sum_{d=0}^{\infty}\frac{\psi_{i,d}}{z^{d}}\in{\mathbb{K}}[[z^{-1}]]^{\times}\right)_{i\in I}$
which are rational, i.e. each $\psi_{i}(z)$ is the power series expansion of a rational function in $z$. Assume that the numerators and denominators of all such $\ell$-weights completely factor into $1-\gamma z^{-1}$ for various $\gamma\in{\mathbb{K}}^{*}$, which is very close to the standard treatment of these objects: typically, $\mathbb{K}$} is replaced by $\mathbb{C}$} and $q$ is replaced by a generic complex number. The modules $L({\boldsymbol{\psi}})$ are completely determined by their $q$-characters ([19])
(2) $\chi_{q}(V)=\sum_{\ell\text{-weights }{\boldsymbol{\psi}}}\dim_{{\mathbb{K}}}(V_{{\boldsymbol{\psi}}})[{\boldsymbol{\psi}}]$
Above, $V_{\boldsymbol{\psi}}$ denotes the generalized eigenspace of $V$ in which the commuting family of operators $\{\varphi_{i,d}^{+}\}_{i\in I,d\geq 0}$ acts according to the eigenvalues $\{\psi_{i,d}\}_{i\in I,d\geq 0}$, and $\boldsymbol{\psi}$}] denotes a certain formal symbol that satisfies ${\boldsymbol{\psi}}][{\boldsymbol{\psi}}^{\prime}]=[{\boldsymbol{\psi}}{\boldsymbol{\psi}}^{\prime}]$ with respect to component-wise multiplication of $I$-tuples of power series. When $\mathfrak{g}$} is of finite type, the computation of $\chi_{q}(L({\boldsymbol{\psi}}))$ and its rich connections with various parts of representation theory, algebraic geometry, mathematical physics and combinatorics has been the study of much recent work ([3, 7, 17, 18, 20, 21, 23, 26, 30, 31, 35, 44]).
1.2. Shuffle algebras
The set $\mathbb{N}$} is assumed to contain 0 throughout the present paper. Shuffle algebras have been known to provide explicit models of $U_{q}(L{\mathfrak{g}})$ since the work of [14, 15]. In more detail, there exists a subspace
${\mathcal{S}}=\bigoplus_{{\boldsymbol{n}}\in{\mathbb{N}^{I}}}{\mathcal{S}}_{{\boldsymbol{n}}}\subseteq\bigoplus_{{\boldsymbol{n}}\in{\mathbb{N}^{I}}}{\mathcal{V}}_{{\boldsymbol{n}}}={\mathcal{V}},\quad\text{where }{\mathcal{V}}_{{\boldsymbol{n}}}={\mathbb{K}}[z_{i1}^{\pm 1},\dots,z_{in_{i}}^{\pm 1}]^{\textrm{sym}}_{i\in I}$
such that $\mathcal{S}$} is isomorphic to the half subalgebra $U_{q}^{-}(L{\mathfrak{g}})\subset U_{q}(L{\mathfrak{g}})$ generated by Drinfeld’s elements $\{f_{i,d}\}_{i\in I,d\in{\mathbb{Z}}}$. In [38], we proved that the simple module $L({\boldsymbol{\psi}})$ factors (as a graded vector space) as
(3) $L({\boldsymbol{\psi}})\cong L^{\text{ord }{\boldsymbol{\psi}}}\otimes L^{\neq 0}({\boldsymbol{\psi}})$
where the graded vector space $L^{\text{ord }{\boldsymbol{\psi}}}$ only depends on the orders of the rational functions $\psi_{i}(z)$ at $z=0$, and was described in [39]. Meanwhile, the vector space $L^{\neq 0}({\boldsymbol{\psi}})$ will be shown in [27] to match the underlying graded vector space of the same-named simple module ([24]) of an appropriately shifted quantum loop algebra. We proved in [38] that this vector space has the following shuffle algebra description
(4) $L^{\neq 0}({\boldsymbol{\psi}})={\mathcal{S}}\Big/\left({\mathcal{S}}\cap\Theta({\boldsymbol{\psi}})\right)$
where $\Theta({\boldsymbol{\psi}})\subseteq{\mathcal{V}}$ is an explicit set of Laurent polynomials that we recall in Subsection 3.6. Moreover, the objects above admit gradings that will be denoted by
(5) $L^{\neq 0}({\boldsymbol{\psi}})=\mathop{\bigoplus_{{\boldsymbol{n}}=(n_{i}\geq 0)_{i\in I}}}_{\boldsymbol{x}=(x_{ia}\in{\mathbb{K}}^{*})^{i\in I}_{1\leq a\leq n_{i}}}{\mathcal{S}}_{{\boldsymbol{n}}}\Big/\left({\mathcal{S}}_{{\boldsymbol{n}}}\cap\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}}\right)$
where we recall $\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}}\subseteq{\mathcal{V}}_{{\boldsymbol{n}}}$ in Subsection 3.6. If $A_{i,x}^{-1}=\left[\left(\frac{z-xq^{d_{ij}}}{zq^{d_{ij}}-x}\right)_{j\in I}\right]$, then
(6) $\chi_{q}(L^{\neq 0}({\boldsymbol{\psi}}))=[{\boldsymbol{\psi}}]\sum_{{\boldsymbol{n}},\boldsymbol{x}\text{ as above}}\dim_{{\mathbb{K}}}\left({\mathcal{S}}_{{\boldsymbol{n}}}\Big/\left({\mathcal{S}}_{{\boldsymbol{n}}}\cap\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}}\right)\right)\prod_{i\in I}\prod_{a=1}^{n_{i}}A_{i,x_{ia}}^{-1}$
see [39, Subsection 4.8]. Formula (3) implies the equality
$\chi_{q}(L({\boldsymbol{\psi}}))=\chi^{\text{ord }{\boldsymbol{\psi}}}\cdot\chi_{q}(L^{\neq 0}({\boldsymbol{\psi}}))$
The factor $\chi^{\text{ord }{\boldsymbol{\psi}}}$ was calculated in [39] (in finite types, in accordance with a conjecture of [31, 49]). Thus, the computation of $q$-characters boils down to understanding the dimensions of the vector spaces ${\mathcal{S}}_{{\boldsymbol{n}}}/\left({\mathcal{S}}_{{\boldsymbol{n}}}\cap\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}}\right)$. For $\mathfrak{g}$} of classical types, these dimensions can in principle be calculated following the residue description of [38], and we expect one may recover the combinatorial formulas of [29, 32, 33, 34, 46].
1.3. Euler characteristics
The main purpose of the present paper is to produce a geometric realization of ${\mathcal{S}}_{{\boldsymbol{n}}}/\left({\mathcal{S}}_{{\boldsymbol{n}}}\cap\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}}\right)$, inspired by the following seminal construction of Hernandez-Leclerc. We assume that  [^1] As has long been known to experts, working with $\boldsymbol{\psi}$} of the form (7) does not represent any actual restriction, since the $q$-character of an arbitrary $L({\boldsymbol{\psi}})$ naturally splits as a product of $q$-characters of $L({\boldsymbol{\psi}})$’s of the form (7): in other words, poles and zeroes of $\boldsymbol{\psi}$} that are in different geometric progressions with ratio $q$ do not interact with each other.
(7) ${\boldsymbol{\psi}}=\left(\psi_{i}(z)=\text{constant}_{i}\cdot\frac{\prod_{s\in{\mathbb{Z}}}\left(1-\frac{q^{s}}{z}\right)^{l_{i,s}}}{\prod_{s\in{\mathbb{Z}}}\left(1-\frac{q^{s}}{z}\right)^{m_{i,s}}}\right)_{i\in I}$
for various $l_{i,s},m_{i,s}\in{\mathbb{N}}$, almost all of which are 0. We will encode these numbers in the vectors of non-negative integers ${\mathbf{l}}=(l_{i,s})_{i\in I,s\in{\mathbb{Z}}}$ and ${\boldsymbol{m}}=(m_{i,s})_{i\in I,s\in{\mathbb{Z}}}$.
Theorem 1.4.
(general conjecture in [26], proof for Kirillov-Reshetikhin $\boldsymbol{\psi}$} in loc. cit., general proof in [30]) For $\mathfrak{g}$} of finite type and “reachable” highest $\ell$-weights $\boldsymbol{\psi}$},
(8) $\chi_{q}(L({\boldsymbol{\psi}}))=[{\boldsymbol{\psi}}]\mathop{\bigoplus_{{\boldsymbol{n}}=(n_{i}\geq 0)_{i\in I}}}_{\boldsymbol{x}=(x_{ia}\in q^{{\mathbb{Z}}})^{i\in I}_{1\leq a\leq n_{i}}}(\text{Euler characteristic of $N^{\emph{stab}}_{\boldsymbol{x},{\boldsymbol{m}},{\mathbf{l}}}$})\prod_{i\in I}\prod_{a=1}^{n_{i}}A_{i,x_{ia}}^{-1}$
where the algebraic variety $N^{\emph{stab}}_{\boldsymbol{x},{\boldsymbol{m}},{\mathbf{l}}}$ is defined as follows:
  • consider the quiver $Q^{\emph{gr}}$ with vertex set $I\times{\mathbb{Z}}$ and arrows $(i,s)\mapsto(j,s-d_{ij})$;
  • consider the moduli space $M^{\emph{stab}}_{\boldsymbol{x},{\boldsymbol{m}}}$ of stable $Q^{\emph{gr}}$ representations of dimension $|\{1\leq a\leq n_{i}\text{ s.t. }x_{ia}=q^{s}\}|$ and framing $m_{i,s}$ at any vertex $(i,s)$ (see Section 2);
  • consider the closed subvariety
    (9) $N^{\emph{stab}}_{\boldsymbol{x},{\boldsymbol{m}},{\mathbf{l}}}\subset M^{\emph{stab}}_{\boldsymbol{x},{\boldsymbol{m}}}$
    defined by intersecting the critical locus of the potential (defined below)
    (10) $\emph{Crit}\left(\emph{Tr}(W^{\emph{gr}})\right)$
    with the closed subvarieties defined by the vanishing condition of $l_{i,s}$ generic linear combinations of framed paths ending at any vertex $(i,s)\in I\times{\mathbb{Z}}$ (see Section 2).
The potential $W^{\emph{gr}}$ in (10) is the sum over all cycles of the form ($\forall i\neq j,\forall s\in{\mathbb{Z}}$)
${(j,s)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(i,s+d_{ij})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(i,s+d_{ij}+d_{ii})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle{\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(i,s-d_{ij}-d_{ii})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(i,s-d_{ij})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
and its trace naturally induces a function on $M^{\emph{stab}}_{\boldsymbol{x},{\boldsymbol{m}}}$. The critical locus of this function, which appears in (10), is explicitly determined by the equations
(11) $\left\{\frac{\partial W^{\emph{gr}}}{\partial e}=0\right\}_{\forall\text{ arrow }e}$
Kashiwara-Kim-Oh-Park proved the result above by establishing a monoidal categorification of the cluster algebra associated to the quiver $Q^{\text{gr}}$ (for which the relation to Euler characteristics of quiver moduli had been established on quite general grounds in [10]), following the expectation of Hernandez-Leclerc. This explains the need for the adjective “reachable” in the statement of Theprem 1.4, since these are precisely the modules whose $q$-characters correspond to cluster monomials.
1.5. General highest $\ell$-weights $\boldsymbol{\psi}$}
We wish to generalize (8) to all highest $\ell$-weights $\boldsymbol{\psi}$}, even those which produce simple modules only of the Borel subalgebra of $U_{q}(L{\mathfrak{g}})$. Firstly, in this generality it is natural to replace $\chi_{q}(L({\boldsymbol{\psi}}))$ with $\chi_{q}(L^{\neq 0}({\boldsymbol{\psi}}))$ in the left-hand side (these two quantities are equal for highest $\ell$-weights for which the action extends to the entire $U_{q}(L{\mathfrak{g}})$). Secondly, it is naive to hope that the right-hand side should be something as nice as the Euler characteristic. Therefore, our intended generalization will be to find a geometric formula for the coefficients of $\chi_{q}(L^{\neq 0}({\boldsymbol{\psi}}))$, in terms of the algebraic variety $N^{\text{stab}}_{\boldsymbol{x},{\boldsymbol{m}},{\mathbf{l}}}$. Since we already know from (6) that the aforementioned coefficients are given by the dimensions of the vector spaces ${\mathcal{S}}_{{\boldsymbol{n}}}/({\mathcal{S}}_{{\boldsymbol{n}}}\cap\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}})$, then the problem becomes equivalent to the following:
[…]
To set up our solution to the problem above, in what follows we fix arbitrary
${\boldsymbol{n}}=(n_{i}\geq 0)_{i\in I}\qquad\text{and}\qquad\boldsymbol{x}=(x_{ia}\in q^{{\mathbb{Z}}})_{i\in I,1\leq a\leq n_{i}}$
Let $K(X)$ be the 0-th (topological or algebraic) $K$-theory ring of a variety $X$.
Theorem 1.6.
For any symmetrizable Kac-Moody $\mathfrak{g}$}, we have a ring isomorphism
(13) $\Gamma:K(M_{\boldsymbol{x},{\boldsymbol{m}}}^{\emph{stab}})\otimes_{{\mathbb{Q}}}{\mathbb{K}}\xrightarrow{\sim}{\mathcal{V}}_{{\boldsymbol{n}}}\Big/\Theta\left(\frac{1}{{\boldsymbol{\psi}}^{\emph{den}}}\right)_{\boldsymbol{x}}$
for any ${\boldsymbol{n}},\boldsymbol{x},{\boldsymbol{m}}$ as above, where ${\boldsymbol{\psi}}^{\emph{den}}$ denotes the denominator of (7). The space $M_{\boldsymbol{x},{\boldsymbol{m}}}^{\emph{stab}}$ has no odd $K$-theory.
Theorem 1.6 is an immediate consequence of Proposition 3.9 and Theorem 2.12. The latter theorem will be proved in Section 2 for moduli spaces of stable framed representations of arbitrary quivers; it complements known descriptions of the cohomology of such moduli spaces from [13] and [16]. Let us set
(14) $\Delta_{{\mathbf{l}}}=\prod_{i\in I}\prod_{a=1}^{n_{i}}\prod_{s\in{\mathbb{Z}}}\left(1-\frac{q^{s}}{z_{ia}}\right)^{l_{i,s}}\in{\mathcal{V}}_{{\boldsymbol{n}}}$
It is easy to show that multiplication by $\Delta_{{\mathbf{l}}}$ gives an isomorphism
(15) ${\mathcal{V}}_{{\boldsymbol{n}}}\Big/\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}}\xrightarrow{\sim}(\Delta_{{\mathbf{l}}}\cdot{\mathcal{V}}_{{\boldsymbol{n}}})\Big/\left((\Delta_{{\mathbf{l}}}\cdot{\mathcal{V}}_{{\boldsymbol{n}}})\cap\Theta\left(\frac{1}{{\boldsymbol{\psi}}^{\text{den}}}\right)_{\boldsymbol{x}}\right)\subseteq{\mathcal{V}}_{{\boldsymbol{n}}}\Big/\Theta\left(\frac{1}{{\boldsymbol{\psi}}^{\text{den}}}\right)_{\boldsymbol{x}}$
which clearly sends
(16) ${\mathcal{S}}_{{\boldsymbol{n}}}\Big/\left({\mathcal{S}}_{{\boldsymbol{n}}}\cap\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}}\right)\xrightarrow{\sim}(\Delta_{{\mathbf{l}}}\cdot{\mathcal{S}}_{{\boldsymbol{n}}})\Big/\left((\Delta_{{\mathbf{l}}}\cdot{\mathcal{S}}_{{\boldsymbol{n}}})\cap\Theta\left(\frac{1}{{\boldsymbol{\psi}}^{\text{den}}}\right)_{\boldsymbol{x}}\right)$
Therefore, the isomorphism (13) allows us to more precisely state problem (12) as
(17) $\textbf{realize }\Gamma^{-1}\left((\Delta_{{\mathbf{l}}}\cdot{\mathcal{S}}_{{\boldsymbol{n}}})\Big/\left((\Delta_{{\mathbf{l}}}\cdot{\mathcal{S}}_{{\boldsymbol{n}}})\cap\Theta\left(\frac{1}{{\boldsymbol{\psi}}^{\text{den}}}\right)_{\boldsymbol{x}}\right)\right)\\ \textbf{in terms of the geometry of }N_{\boldsymbol{x},{\boldsymbol{m}},{\mathbf{l}}}^{\text{stab}}\subset M_{\boldsymbol{x},{\boldsymbol{m}}}^{\text{stab}}$
For example, it is quite conceivable that for the reachable highest $\ell$-weights $\boldsymbol{\psi}$} to which Theorem 1.4 applies, the direct image map
(18) $K(N_{\boldsymbol{x},{\boldsymbol{m}},{\mathbf{l}}}^{\text{stab}})\rightarrow K(M_{\boldsymbol{x},{\boldsymbol{m}}}^{\text{stab}})$
is injective and its image coincides with the vector space that appears on the first line of (17). If $N_{\boldsymbol{x},{\boldsymbol{m}},{\mathbf{l}}}^{\text{stab}}$ could also be shown to not have any odd $K$-theory, then the previous sentence would provide an alternate proof of Theorem 1.4.
1.7. Critical $K$-theory
Let us consider critical $K$-theory, which was developed in [1, 12, 28] (and more recently [6]) and applied to the understanding of certain simple modules $L({\boldsymbol{\psi}})$ in [47, 48]; see also [43] for the original definition of $K$-theoretic Hall algebras. We will start from a certain quiver with potential which has been studied in [22]: let $Q$ have vertex set $I$ and a single arrow $_{j}\square_{i}:i\rightarrow j$ for all $i,j\in I$ (this includes a single loop at every vertex $i$). Then we consider the potential
(19) $W=\sum_{i\neq j}{{}_{j}\square_{i}}\underbrace{{{}_{i}\square_{i}}\dots{}_{i}\square_{i}}_{-c_{ij}\text{ factors}}{{}_{i}\square_{j}}$
[…]
$K^{{\mathbb{C}}^{*}}(Q,W)_{{\boldsymbol{n}}}\xrightarrow{\iota_{{\boldsymbol{n}}}}K^{{\mathbb{C}}^{*}}(M_{{\boldsymbol{n}}})={\mathcal{V}}_{{\boldsymbol{n}}}$
The map $\iota_{{\boldsymbol{n}}}$ is a ${\mathcal{V}}_{{\boldsymbol{n}}}$-module homomorphism with respect to natural module structures on the domain and target. Thus, there is a well-defined notion of multiplication on $K^{{\mathbb{C}}^{*}}(Q,W)_{{\boldsymbol{n}}}$ by $\Delta_{{\mathbf{l}}}$ of (14). In Section 4, we will encounter the diagram
(20) ${K^{{\mathbb{C}}^{*}}(Q,W)_{{\boldsymbol{n}}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$$\scriptstyle{\cdot\Delta_{{\mathbf{l}}}}$${K^{{\mathbb{C}}^{*}}(Q,W)_{{\boldsymbol{n}}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\iota_{{\boldsymbol{n}}}}$${K^{{\mathbb{C}}^{*}}(M_{{\boldsymbol{n}}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${K(Q^{\text{gr}},W^{\text{gr}})^{\text{stab}}_{\boldsymbol{x},{\boldsymbol{m}}}\otimes_{{\mathbb{Q}}}{\mathbb{K}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\cdot\Delta^{\text{gr}}_{{\mathbf{l}}}}$${K(Q^{\text{gr}},W^{\text{gr}})^{\text{stab}}_{\boldsymbol{x},{\boldsymbol{m}}}\otimes_{{\mathbb{Q}}}{\mathbb{K}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\iota^{\text{gr},\text{stab}}_{\boldsymbol{x},{\boldsymbol{m}}}}$${K(M_{\boldsymbol{x},{\boldsymbol{m}}}^{\text{stab}})\otimes_{{\mathbb{Q}}}{\mathbb{K}}}$
where the spaces in the bottom row are analogues of those in the top row for the quiver of Theorem 1.4, intersected with the stable open locus (see Subsection 4.6 for more details). In the bottom row, $\Delta_{{\mathbf{l}}}^{\text{gr}}$ is (a slightly renormalized version of) the product of those linear factors of (14) which vanish at $z_{ia}=x_{ia}$, specifically
(21) $\Delta_{{\mathbf{l}}}^{\text{gr}}=\prod_{i\in I}\prod_{a=1}^{n_{i}}\underbrace{\left(1-\frac{1}{z_{ia}}\right)^{l_{i,s}}}_{s\text{ is defined by }x_{ia}=q^{s}}$
The map $\pi$ in (20) is given by restriction, while $\iota_{{\boldsymbol{n}}},\iota^{\text{gr},\text{stab}}_{\boldsymbol{x},{\boldsymbol{m}}}$ behave like push-forwards, and so diagram (20) does not commute on the nose. However, we will show that
[…]
Theorem 1.8.
For any strongly symmetrizable  [^2] See (65) for the definition of strongly symmetrizable Kac-Moody Lie algebras. We note that all finite type and affine type Lie algebras are strongly symmetrizable, except for $\widehat{{\mathfrak{sl}}}_{2}$. Kac-Moody Lie algebra $\mathfrak{g}$} and any highest $\ell$-weight $\boldsymbol{\psi}$}, the image of the composition $\rightarrow\rightarrow\downarrow$ in diagram (20) is naturally isomorphic to the vector space
(23) ${\mathcal{S}}_{{\boldsymbol{n}}}\Big/\left({\mathcal{S}}_{{\boldsymbol{n}}}\cap\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}}\right)$
and thus has dimension equal to the appropriate coefficient of $\chi_{q}(L^{\neq 0}({\boldsymbol{\psi}}))$.
We conjectured the following result in the first draft of the present paper, and were later informed by Cao-Okounkov-Zhou-Zhou that it follows from their general theory of stable envelopes in critical $K$-theory ([5]).
Theorem 1.9.
(conjectured by the author, proved by [5]) The maps $\pi$ and $\iota^{\emph{gr},\emph{stab}}_{\boldsymbol{x},{\boldsymbol{m}}}$ in (20) are surjective and injective, respectively.
Combining Theorems 1.8 and 1.9 with statement (22) implies the following.
Corollary 1.10.
For strongly symmetrizable $\mathfrak{g}$} and any highest $\ell$-weight $\boldsymbol{\psi}$}, we have
(24) ${\mathcal{S}}_{{\boldsymbol{n}}}\Big/\left({\mathcal{S}}_{{\boldsymbol{n}}}\cap\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}}\right)\cong\left[\Delta^{\emph{gr}}_{{\mathbf{l}}}\cdot K(Q^{\emph{gr}},W^{\emph{gr}})^{\emph{stab}}_{{\boldsymbol{m}},\boldsymbol{x}}\right]\otimes_{{\mathbb{Q}}}{\mathbb{K}}$
We claim that the isomorphism (24) provides a solution to problems (12) and (17), and thus generalizes and categorifies Theorem 1.4 (using the fact that critical $K$-theory is categorified by derived factorization categories). To see this, note that the critical $K$-theory in the right-hand side of (24) is a replacement of the $K$-theory of the critical locus of $W^{\text{gr}}$, while multiplication by $\Delta^{\text{gr}}_{{\mathbf{l}}}$ is a derived version of imposing the vanishing of the $l_{i,s}$ linear combinations of framed paths that cut out the closed subvariety (9). Thus, the right-hand side of (24) is a replacement of $K(N_{\boldsymbol{x},{\boldsymbol{m}},{\mathbf{l}}}^{\text{stab}})\otimes_{{\mathbb{Q}}}{\mathbb{K}}$, which in turn categorifies the Euler characteristic in (8). Meanwhile, the left-hand side of (24) categorifies the coefficients of the $q$-character of $L({\boldsymbol{\psi}})$, due to (5).
1.11. Conclusion
To summarize, we provide a geometric construction of all simple modules $L({\boldsymbol{\psi}})$ in the Borel category $\mathcal{O}$} of quantum loop algebras $U_{q}(L{\mathfrak{g}})$, for any strongly symmetrizable Kac-Moody Lie algebra $\mathfrak{g}$} (which includes all finite and affine $\mathfrak{g}$} except for $\widehat{{\mathfrak{sl}}}_{2}$). This fits into a very active field of research, that started with the Euler characteristic construction of [26], continued with the connections with physics explored in [4], and more recently involved the construction of Kirillov-Reshetikhin and prefundamental modules via critical $K$-theory in [47, 48]. While we consider similar objects as the latter two papers, our approach is somewhat different, and in a certain sense complementary to theirs. To explain this last comment, recall that in loc. cit. the authors defined a convolution product on
(25) $K^{{\mathbb{C}}^{*}}(Q^{\text{gr}},W^{\text{gr}})=\bigoplus_{{\boldsymbol{n}}\in{\mathbb{N}^{I}}}K^{{\mathbb{C}}^{*}}(Q^{\text{gr}},W^{\text{gr}})_{{\boldsymbol{n}}}$
related it to the quantum affine algebra $U_{q}(L{\mathfrak{g}})$, and defined prefundamental modules (i.e. the case ${\mathbf{l}}={\boldsymbol{0}}$ in (7)) using critical $K$-theory for the same quiver with potential as us. On the other hand, our viewpoint is to start from the vector space
(26) $\bigoplus_{{\boldsymbol{n}},\boldsymbol{x}}{\mathcal{S}}_{{\boldsymbol{n}}}\Big/\left({\mathcal{S}}_{{\boldsymbol{n}}}\cap\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}}\right)$
which we already know from [38] yields the simple module $L({\boldsymbol{\psi}})$, and then seek geometric realizations for the direct summands in (26). With this in mind, our Corollary 1.10 not only provides a geometric construction of all simple modules of quantum affine algebras, but provides an explicit computation of the corresponding critical $K$-theory in terms of shuffle algebras. We note that Proposition 4.2 also provides a shuffle algebra computation of the critical K-HA studied in [48].
1.12. Acknowledgements
I would like to thank Ben Davison, David Hernandez, Shivang Jindal, Hiraku Nakajima, Tudor Pădurariu, Markus Reineke and Yukinobu Toda for numerous important conversations on the topics of the present paper, and the authors of [5] for their proof of Theorem 1.9.
2. Moduli of quiver representations
We recall the cell decomposition of moduli spaces of stable framed quiver representations from [13], and use it to describe the ring structure on the $K$-theory of the aforementioned moduli spaces (in a different way from the earlier work [16]). Hereafter, we write $K(X)$ for the 0-th $K$-theory  [^3] The contents of the present Section hold equally well for topological and algebraic $K$-theory; the former is closer in spirit to [26], while the latter is closer technically to Section 4. ring of a smooth complex algebraic variety $X$, with rational coefficients. We thank Markus Reineke for sharing with us much advice pertaining to the contents of the present Section.
2.1. Framed quiver representations
Consider an arbitrary quiver (i.e. directed graph) $Q$ with vertex set $I$ and arrow set $E$. For any ${\boldsymbol{n}}\in{\mathbb{N}^{I}}$, the moduli stack of $\boldsymbol{n}$}-dimensional quiver representations is
(27) $M_{{\boldsymbol{n}}}=\left(\bigoplus_{e=\overrightarrow{ij}\in E}\text{Hom}({\mathbb{C}}^{n_{i}},{\mathbb{C}}^{n_{j}})\right)\Big/\prod_{i\in I}GL_{n_{i}}$
where automorphisms $(g_{i})_{i\in I}$ act on representations $(\phi_{e}:{\mathbb{C}}^{n_{\text{tail}(e)}}\rightarrow{\mathbb{C}}^{n_{\text{head}(e)}})_{e\in E}$ by conjugation. Given any ${\boldsymbol{m}}\in{\mathbb{N}^{I}}$, we may also consider the moduli stack of framed $\boldsymbol{n}$}-dimensional quiver representations
(28) $M_{{\boldsymbol{n}},{\boldsymbol{m}}}=\left(\bigoplus_{i\in I}\text{Hom}({\mathbb{C}}^{m_{i}},{\mathbb{C}}^{n_{i}})\bigoplus_{e=\overrightarrow{ij}\in E}\text{Hom}({\mathbb{C}}^{n_{i}},{\mathbb{C}}^{n_{j}})\right)\Big/\prod_{i\in I}GL_{n_{i}}$
Thus, a framed quiver representation consists of linear maps
(29) $\Big(\psi_{i}:{\mathbb{C}}^{m_{i}}\rightarrow T_{i},\ \phi_{e}:T_{\text{tail}(e)}\rightarrow T_{\text{head}(e)}\Big)_{\forall i\in I,e\in E}$
where $T_{i}$ is a $n_{i}$-dimensional vector space. Such a framed quiver representation is called stable if it has no proper subrepresentation which contains the images of all the $\psi_{i}$’s. The open substack of stable framed quiver representations
(30) $M^{\text{stab}}_{{\boldsymbol{n}},{\boldsymbol{m}}}=\left(\bigoplus_{i\in I}\text{Hom}({\mathbb{C}}^{m_{i}},{\mathbb{C}}^{n_{i}})\bigoplus_{e=\overrightarrow{ij}\in E}\text{Hom}({\mathbb{C}}^{n_{i}},{\mathbb{C}}^{n_{j}})\right)^{\text{stable}}\Big/\prod_{i\in I}GL_{n_{i}}$
is actually a smooth variety. Let us consider the torus action
(31) ${\mathbb{C}}^{*}\curvearrowright M^{\text{stab}}_{{\boldsymbol{n}},{\boldsymbol{m}}}$
which scales all the linear maps $\phi_{e}$ in (29) with weigth 1. The fixed point set
(32) $M^{\text{stab},\text{nilp}}_{{\boldsymbol{n}},{\boldsymbol{m}}}=\left(M^{\text{stab}}_{{\boldsymbol{n}},{\boldsymbol{m}}}\right)^{{\mathbb{C}}^{*}}$
is well-known to be proper and isomorphic to the moduli space of stable framed nilpotent quiver representations, i.e. the set of collections (29) for which there exists $N\gg 0$ such that the composition of any $N$ of the maps $\phi_{e}$ is 0.
2.2. Paths
We will fix $\boldsymbol{m}$} and $\boldsymbol{n}$} throughout the present section. The moduli spaces of stable framed quiver representations were studied in [13], where a cell decomposition that we now recall was first constructed.
Definition 2.3.
Whereas a path in the quiver $Q$ is a chain of arrows $e_{k}\dots e_{1}$ with $\emph{tail}(e_{s+1})=\emph{head}(e_{s})$ for all $s\in\{1,\dots,k-1\}$, we refer to a framed path as a chain
(33) $p=e_{k}\dots e_{1}o_{i,\ell}$
for some $\ell\in\{1,\dots,m_{i}\}$, where $i=\emph{tail}(e_{1})$. Intuitively, $o_{i,\ell}$ is a symbol that corresponds to the $\ell$-th coordinate of the linear map $\psi_{i}$ in (29). We will call
$\emph{head}(p)=\emph{head}(e_{k})\in I$
the head of a framed path $p$ as in (33).
Just like a path $e_{k}\dots e_{1}$ gives rise to a map $\phi_{e_{k}}\dots\phi_{e_{1}}:T_{\text{tail}(e_{1})}\rightarrow T_{\text{head}(e_{k})}$ in any quiver representation, a framed path $p$ as in (33) gives rise to a vector
$v_{p}=\phi_{e_{k}}\dots\phi_{e_{1}}\psi_{i}(\ell\text{-th standard basis vector of }{\mathbb{C}}^{m_{i}})\in T_{\text{head}(p)}$
in any framed quiver representation (29).
2.4. Cell decompositions
Let us consider a total order on the set of framed paths $p$ as in (33), which has the property that
$p<ep$
for any arrow $e$ whose tail matches the head of $p$. This order induces a lexicographic order on collections of framed paths, namely
(34) $S=\Big\{p_{1}<\dots<p_{n}\Big\}$
We will only consider such collections for $n=\sum_{i\in I}n_{i}$, given our fixed ${\boldsymbol{n}}=(n_{i})_{i\in I}\in{\mathbb{N}^{I}}$. If we let $i_{1},\dots,i_{n}\in I$ denote the heads of $p_{1},\dots,p_{n}$ respectively, then
(35) $\boldsymbol{\varsigma}^{i_{1}}+\dots+\boldsymbol{\varsigma}^{i_{n}}={\boldsymbol{n}}$
where $\boldsymbol{\varsigma}^{i}\in{\mathbb{N}^{I}}$ denotes the $I$-tuple with entry $1$ on position $i$ and 0 everywhere else. Whenever (35) holds, we call $i_{1},\dots,i_{n}$ an ordering of $\boldsymbol{n}$}. We will only consider collections $S$ which are acceptable, in the sense that every $p_{b}$ is either equal to $o_{i,\ell}$ (for some $i\in I$, $\ell\in\{1,\dots,w_{i}\}$) or to $ep_{a}$ (with $a<b$ and $e\in E$ having tail equal to $\text{head}(p_{a})$). The following construction is a key result of [13] (see also [16]).
Definition 2.5.
For any acceptable collection $S=\{p_{1}<\dots<p_{n}\}$, consider
(36) $Z_{S}\subset M^{\emph{stab}}_{{\boldsymbol{n}},{\boldsymbol{m}}}$
consisting of those stable framed quiver representations for which $v_{p_{1}},\dots,v_{p_{n}}$ form a basis of the vector spaces $\{T_{i}\}_{i\in I}$, and moreover there exist $\gamma_{a,b,e}\in{\mathbb{C}}$ such that
(37) $\phi_{e}(v_{p_{a}})=\sum_{1\leq b\leq n|ep_{a}>p_{b}}\gamma_{a,b,e}v_{p_{b}}$
for any $a\in\{1,\dots,n\}$ and any $e\in E$ such that $\emph{tail}(e)=\emph{head}(p_{a})$ but $ep_{a}\notin S$.
It was shown in [13] that the coefficients $\gamma_{a,b,e}$ in (37) give rise to an isomorphism
(38) $Z_{S}\cong{\mathbb{A}}^{\#\{1\leq a,b\leq n,e\in E|\text{tail}(e)=\text{head}(p_{a}),ep_{a}>p_{b},ep_{a}\notin S\}}$
and that we have a cell decomposition
(39) $M^{\text{stab}}_{{\boldsymbol{n}},{\boldsymbol{m}}}=\bigsqcup_{\text{acceptable collections }S}Z_{S}$
It is well-known that the existence of the cell decomposition (39) implies that the 0-th $K$-theory (with $\mathbb{Q}$}-coefficients, throughout this Section) of $M^{\text{stab}}_{{\boldsymbol{n}},{\boldsymbol{m}}}$ satisfies
(40) $\dim_{{\mathbb{Q}}}\left(K(M^{\text{stab}}_{{\boldsymbol{n}},{\boldsymbol{m}}})\right)=\Big|\Big\{\text{acceptable collections}\Big\}\Big|$
and there is no odd $K$-theory. Recall the torus action (31), whose fixed point set is the moduli space of nilpotent stable framed quiver representations.
Proposition 2.6.
If we let $Z_{S}^{\emph{nilp}}=Z_{S}\cap M_{{\boldsymbol{m}},{\boldsymbol{n}}}^{\emph{stab},\emph{nilp}}$, then
(41) $M_{{\boldsymbol{n}},{\boldsymbol{m}}}^{\emph{stab},\emph{nilp}}=\bigsqcup_{\text{acceptable collections }S}Z^{\emph{nilp}}_{S}$
is also a cell decomposition, so we conclude that
(42) $\dim_{{\mathbb{Q}}}\left(K(M^{\emph{stab},\emph{nilp}}_{{\boldsymbol{n}},{\boldsymbol{m}}})\right)=\Big|\Big\{\emph{acceptable collections}\Big\}\Big|$
and $M^{\emph{stab},\emph{nilp}}_{{\boldsymbol{n}},{\boldsymbol{m}}}$ has no odd $K$-theory.
Proof.
Recall that $Z_{S}^{\text{nilp}}=Z_{S}^{{\mathbb{C}}^{*}}$ with respect to the $\mathbb{C}$$^{*}$ action that scales all the arrows with weight 1. With respect to the isomorphism (38), the induced action
${\mathbb{C}}^{*}\curvearrowright{\mathbb{A}}^{\#\{1\leq a,b\leq n,e\in E|\text{tail}(e)=\text{head}(p_{a}),ep_{a}>p_{b},ep_{a}\notin S\}}$
simply scales each coordinate $\gamma_{a,b,e}$ by a certain power of the tautological character of $\mathbb{C}$$^{*}$ (the power in question is 1 plus the number of consituent arrows of the path $p_{a}$ minus the number of constituent arrows of the path $p_{b}$). Therefore, $Z_{S}^{\text{nilp}}$ is isomorphic to the affine subspace determined by those coordinates for which the aforementioned power is 0. Then (42) follows from (41) as (40) followed from (39).
2.7. Interlude on quiver Grassmannians
The moduli space $M^{\text{stab}}_{{\boldsymbol{n}},{\boldsymbol{m}}}$ has an alternative interpretation in terms of modules over the path algebra
${\mathbb{C}}Q=\bigoplus_{\text{path }p=e_{k}\dots e_{1}}{\mathbb{C}}p$
with product given by concatenation of paths (we consider the empty paths $\varnothing_{i}$ at each $i\in I$ to be distinct idempotents in the path algebra). There is a one-to-one correspondence between representations of the quiver $Q$ and ${\mathbb{C}}Q$-modules. Let us consider the dual injective ${\mathbb{C}}Q$-module with socle $i\in I$
$I_{i}^{*}=\bigoplus_{\text{path }p=e_{k}\dots e_{1}\text{ starting at }i}{\mathbb{C}}p$
with the action via concatenation. By a slight abuse of notation, we will write $I_{i}^{*}$ for the corresponding quiver representation, and note that it is infinite-dimensional if $Q$ has oriented cycles. In the following definition, fix a quiver representation
$T=\{\phi_{e}:T_{i}\rightarrow T_{j}\}_{e=\overrightarrow{ij}\in E}$
Definition 2.8.
For any ${\boldsymbol{n}}\in{\mathbb{N}^{I}}$, the quiver Grassmannian is
(43) $\emph{Gr}_{{\boldsymbol{n}}}(T)=\Big\{\left(T_{i}\stackrel{{\scriptstyle\pi_{i}}}{{\twoheadrightarrow}}{\mathbb{C}}^{n_{i}}\right)_{i\in I}\text{ compatibly with the }\phi_{e}\text{'s}\Big\}\Big/\prod_{i\in I}GL_{n_{i}}$
where the compatibility condition means that every $\phi_{e}$ takes $\emph{Ker }\pi_{i}$ to $\emph{Ker }\pi_{j}$ (in other words, the quotients induce a quiver representation structure on $\{{\mathbb{C}}^{n_{i}}\}_{i\in I}$).
When $T$ is finite-dimensional, it is clear that $\text{Gr}_{{\boldsymbol{n}}}(T)$ is a closed subset of a product of ordinary Grassmannians, and thus itself a projective variety. However, we will actually be most interested in the following setup.
Proposition 2.9.
For any ${\boldsymbol{m}},{\boldsymbol{n}}\in{\mathbb{N}^{I}}$, we have an isomorphism
(44) $\emph{Gr}_{{\boldsymbol{n}}}\left(\bigoplus_{k\in I}(I_{k}^{*})^{\oplus m_{k}}\right)\cong M^{\emph{stab}}_{{\boldsymbol{n}},{\boldsymbol{m}}}$
Proof.
Recall that $I_{k}^{*}$ is freely generated by the action of ${\mathbb{C}}Q$ on the empty path $\varnothing_{k}$ at $k$, hence $(I_{k}^{*})_{i}$ has a basis indexed by paths $p:k\rightarrow i$. Thus, to give surjections
(45) $\left(\bigoplus_{k\in I}\underbrace{(I_{k}^{*})_{i}\oplus\dots\oplus(I_{k}^{*})_{i}}_{m_{k}\text{ summands}}\twoheadrightarrow{\mathbb{C}}^{n_{i}}\right)_{i\in I}$
one needs to provide spanning sets
(46) $\Big(v_{p}^{(k,1)},\dots,v_{p}^{(k,m_{k})}\Big)_{\text{paths }p:k\rightarrow i}\quad\text{of}\quad{\mathbb{C}}^{n_{i}}$
for all $i\in I$. In order for this assignment to be compatible in the sense of (43), the induced quiver maps $\phi_{e}:{\mathbb{C}}^{n_{\text{head}(e)}}\rightarrow{\mathbb{C}}^{n_{\text{tail}(e)}}$ must satisfy
$\phi_{e}(v_{p}^{(k,a)})=\begin{cases}v_{ep}^{(k,a)}&\text{if }\text{tail}(p)=\text{head}(e)\\ 0&\text{otherwise}\end{cases}$
for all $k\in I$ and $a\in\{1,\dots,m_{k}\}$. However, the equation above means that the datum (46) is completely determined by the collection of vectors
$\Big(v_{\varnothing_{k}}^{(k,1)},\dots,v_{\varnothing_{k}}^{(k,m_{k})}\Big)\subset{\mathbb{C}}^{n_{k}}$
as $k\in I$ varies, which altogether generate the quiver representation $\{{\mathbb{C}}^{n_{i}}\}_{i\in I}$. In turn, this datum entails the same information as a point of $M^{\text{stab}}_{{\boldsymbol{n}},{\boldsymbol{m}}}$, with $v_{\varnothing_{k}}^{(k,a)}$ being the image of the $a$-th standard basis vector under the framing map ${\mathbb{C}}^{m_{k}}\rightarrow{\mathbb{C}}^{n_{k}}$.
If we let $\text{Gr}^{\text{nilp}}_{{\boldsymbol{n}}}(T)$ denote the closed subset of $\text{Gr}_{{\boldsymbol{n}}}(T)$ such that the induced quiver representation $({\mathbb{C}}^{n_{i}})_{i\in I}$ is nilpotent, then we have
$\text{Gr}^{\text{nilp}}_{{\boldsymbol{n}}}\left(\bigoplus_{k\in I}(I_{k}^{*})^{\oplus m_{k}}\right)\cong M_{{\boldsymbol{n}},{\boldsymbol{m}}}^{\text{stab},\text{nilp}}$
By dualizing the surjective map of quiver representations (45), we see that
(47) $\text{Gr}^{\text{nilp}}_{{\boldsymbol{n}}}\left(\bigoplus_{k\in I}(I_{k}^{*})^{\oplus m_{k}}\right)$ $=\Big\{{\boldsymbol{n}}\text{-dim nilpotent subreps of }\bigoplus_{k\in I}(I_{k}^{**})^{\oplus m_{k}}\Big\}$
$=\Big\{{\boldsymbol{n}}\text{-dim subreps of }\bigoplus_{k\in I}I_{k}^{\oplus m_{k}}\Big\}$
which is the description of the quiver Grassmannian in [10] and [26].
2.10. Residues
While acceptable collections provide a great indexing set for the cells (and thus for a basis of the $K$-theory) of $M_{{\boldsymbol{n}},{\boldsymbol{m}}}^{\text{stab}}$, we will now give a description of the latter $K$-theory’s underlying ring structure. To this end, consider the ring
(48) $V_{{\boldsymbol{n}}}={\mathbb{Q}}[z_{i1}^{\pm 1},\dots,z_{in_{i}}^{\pm 1}]^{\textrm{sym}}_{i\in I}$
where the superscript “sym” refers to Laurent polynomials which are symmetric in the variables $z_{i1},\dots,z_{in_{i}}$ for each $i$ separately (such Laurent polynomials will be called “color-symmetric”). It is well-known that we have an isomorphism
(49) $V_{{\boldsymbol{n}}}\xrightarrow{\sim}K(M_{{\boldsymbol{n}}})$
which sends the $k$-th elementary symmetric function in $z_{i1},\dots,z_{in_{i}}$ to the $k$-th exterior power of the tautological rank $n_{i}$ vector bundle
(50) ${{\mathcal{T}}_{i}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${M_{{\boldsymbol{n}}}}$
which parameterizes the vector spaces ${\mathbb{C}}^{n_{i}}$ in (27) (in what follows, we will also write ${\mathcal{T}}_{i}$ for the pull-back of the above vector bundle to various variants of $M_{{\boldsymbol{n}}}$, such as moduli of stable framed quiver representations). $M_{{\boldsymbol{n}}}$ has no odd $K$-theory.
Definition 2.11.
For any $i_{1},\dots,i_{n}\in I$, consider the set
(51) $\Theta_{i_{1},\dots,i_{n}}^{{\boldsymbol{m}}}\subseteq{\mathbb{Q}}[z_{1}^{\pm 1},\dots,z_{n}^{\pm 1}]$
consisting of Laurent polynomials $F(z_{1},\dots,z_{n})$ such that
(52) $\underset{z_{n}=1}{\emph{Res}}\dots\underset{z_{1}=1}{\emph{Res}}\left(\frac{z_{1}^{d_{1}}\dots z_{n}^{d_{n}}F(z_{1},\dots,z_{n})}{\prod_{a=1}^{n}(z_{a}-1)^{m_{i_{a}}}}\prod_{1\leq a<b\leq n}(z_{b}-z_{a})^{\delta_{i_{a}i_{b}}-|\{\emph{arrows }\overrightarrow{i_{a}i_{b}}\}|}\right)=0$
for all $d_{1},\dots,d_{n}\in{\mathbb{Z}}$.
It is easy to see that $\Theta_{i_{1},\dots,i_{n}}^{{\boldsymbol{m}}}$ is a finite-codimension ideal, since condition (52) boils down to the vanishing of finitely many linear combinations of derivatives of $F$ at the point $(1,\dots,1)$. For any ordering $i_{1},\dots,i_{n}$ of $\boldsymbol{n}$} as in (35), we have an injection
(53) $V_{{\boldsymbol{n}}}\hookrightarrow{\mathbb{Q}}[z_{1}^{\pm 1},\dots,z_{n}^{\pm 1}]$
given by matching each variable $\{z_{a}\}_{1\leq a\leq n}$ with some variable of the form $z_{i_{a}\bullet_{a}}$ in (48) (the choice of numbers $\bullet_{a}$ does not matter, as long as $\bullet_{a}\neq\bullet_{b}$ whenever $a\neq b$, $i_{a}=i_{b}$). This will allow us to intersect the ideals (51) with the ring $V_{{\boldsymbol{n}}}$, which implicitly will mean pulling back $\Theta_{i_{1},\dots,i_{n}}^{{\boldsymbol{m}}}$ under the homomorphism (53).
Theorem 2.12.
The exterior powers of $\{{\mathcal{T}}_{i}\}_{i\in I}$ induce a ring homomorphism
(54) $\rho:V_{{\boldsymbol{n}}}\rightarrow K(M_{{\boldsymbol{n}},{\boldsymbol{m}}}^{\emph{stab}})$
The homomorphism $\rho$ is surjective and its kernel is isomorphic to
(55) $\Theta_{{\boldsymbol{n}}}^{{\boldsymbol{m}}}=\bigcap_{i_{1},\dots,i_{n}\text{ ordering of }{\boldsymbol{n}}}\Theta_{i_{1},\dots,i_{n}}^{{\boldsymbol{m}}}$
Thus, $K\left(M_{{\boldsymbol{n}},{\boldsymbol{m}}}^{\emph{stab}}\right)\cong V_{{\boldsymbol{n}}}/\Theta_{{\boldsymbol{n}}}^{{\boldsymbol{m}}}$. Moreover, $M_{{\boldsymbol{n}},{\boldsymbol{m}}}^{\emph{stab}}$ has no odd $K$-theory.
In particular, Theorem 2.12 and formula (42) imply that
(56) $\dim_{{\mathbb{Q}}}\left(V_{{\boldsymbol{n}}}/\Theta_{{\boldsymbol{n}}}^{{\boldsymbol{m}}}\right)=\Big|\Big\{\text{acceptable collections}\Big\}\Big|$
However, logically we will prove (56) and then use it to deduce Theorem 2.12.
Proof.
For any ordering $i_{1},\dots,i_{n}$ of $\boldsymbol{n}$}, consider the following “flag” version of the framed moduli space (30)
(57) $M^{\text{stab}}_{i_{1},\dots,i_{n},{\boldsymbol{m}}}=$
$\left(\bigoplus_{i\in I}\text{Hom}({\mathbb{C}}^{m_{i}},{\mathbb{C}}^{n_{i}})\bigoplus_{e=\overrightarrow{ij}\in E}\text{Hom}_{i_{1},\dots,i_{n}}({\mathbb{C}}^{n_{i}},{\mathbb{C}}^{n_{j}})\right)^{\text{stable}}\Big/B_{i_{1},\dots,i_{n}}$
where we fix decompositions ${\mathbb{C}}^{n_{i}}=\oplus^{1\leq a\leq n}_{i_{a}=i}{\mathbb{C}}v_{a}$, and we write $\text{Hom}_{i_{1},\dots,i_{n}}({\mathbb{C}}^{n_{i}},{\mathbb{C}}^{n_{j}})$ for the linear space of homomorphisms ${\mathbb{C}}^{n_{i}}\rightarrow{\mathbb{C}}^{n_{j}}$ that take each $v_{a}$ to a linear combination of $v_{b}$ with $a<b$ (similarly, $B_{i_{1},\dots,i_{n}}\subset\prod_{i\in I}GL_{n_{i}}$ refers to the subgroup of automorphisms which take each $v_{a}$ to a linear combination of $v_{b}$ with $a\leq b$). In other words, (57) is the moduli space of stable framed $\boldsymbol{n}$}-dimensional quiver representations which admit a full flag of subrepresentations with successive quotients being the one-dimensional representations supported at vertices $i_{1},\dots,i_{n}\in I$ and zero maps. We have a natural map
(58) $\bigsqcup_{i_{1},\dots,i_{n}\text{ ordering of }{\boldsymbol{n}}}M^{\text{stab}}_{i_{1},\dots,i_{n},{\boldsymbol{m}}}\xrightarrow{\upsilon}M^{\text{stab}}_{{\boldsymbol{n}},{\boldsymbol{m}}}$
whose image is $M^{\text{stab},\text{nilp}}_{{\boldsymbol{n}},{\boldsymbol{m}}}$ (this is obvious from the quiver moduli description: a nilpotent quiver representation is precisely one which can be filtered such that the successive quotients are one-dimensional quiver representations with zero maps).
Lemma 2.13.
We have an isomorphism
(59) $K(M^{\emph{stab}}_{i_{1},\dots,i_{n},{\boldsymbol{m}}})\cong{\mathbb{Q}}[z_{1}^{\pm 1},\dots,z_{n}^{\pm 1}]\Big/\Theta^{{\boldsymbol{m}}}_{i_{1},\dots,i_{n}}$
and $M^{\emph{stab}}_{i_{1},\dots,i_{n},{\boldsymbol{m}}}$ has no odd $K$-theory.
We will prove Lemma 2.13 as soon as we finish the proof of Theorem 2.12. To the latter end, consider the maps
$V_{{\boldsymbol{n}}}\xrightarrow{\rho}K(M_{{\boldsymbol{n}},{\boldsymbol{m}}}^{\text{stab}})\xrightarrow{\upsilon^{*}}\mathop{\bigoplus_{i_{1},\dots,i_{n}}}_{\text{ordering of }{\boldsymbol{n}}}K(M^{\text{stab}}_{i_{1},\dots,i_{n},{\boldsymbol{m}}})\cong\mathop{\bigoplus_{i_{1},\dots,i_{n}}}_{\text{ordering of }{\boldsymbol{n}}}{\mathbb{Q}}[z_{1}^{\pm 1},\dots,z_{n}^{\pm 1}]\Big/\Theta^{{\boldsymbol{m}}}_{i_{1},\dots,i_{n}}$
whose composition corresponds to sending a symmetric polynomial in $\{z_{i1},\dots,z_{in_{i}}\}_{i\in I}$ to the same symmetric polynomial in $\{z_{1},\dots,z_{n}\}$ (with each $z_{a}$ identified with $z_{i_{a}\bullet_{a}}$ for some numbers $\bullet_{a}$, whose particular value does not matter as long as $\bullet_{a}\neq\bullet_{b}$ whenever $a\neq b,i_{a}=i_{b}$). Therefore, we conclude that
$\text{Ker }\rho\subseteq\bigcap_{i_{1},\dots,i_{n}\text{ ordering of }{\boldsymbol{n}}}\Theta^{{\boldsymbol{m}}}_{i_{1},\dots,i_{n}}=\Theta_{{\boldsymbol{n}}}^{{\boldsymbol{m}}}$
and so
$\dim_{{\mathbb{Q}}}\left(K(M^{\text{stab}}_{{\boldsymbol{n}},{\boldsymbol{m}}})\right)\geq\dim_{{\mathbb{Q}}}\left(V_{{\boldsymbol{n}}}/\text{Ker }\rho\right)\geq\dim_{{\mathbb{Q}}}\left(V_{{\boldsymbol{n}}}/\Theta_{{\boldsymbol{n}}}^{{\boldsymbol{m}}}\right)$
Once we prove (56), then formula (42) will imply that all of the inequalities above are actually equalities, and the proof of Theorem 2.12 will be complete.
In order to give a direct proof of (56), let us consider the following linear map for any acceptable collection $S=\{p_{1}<\dots<p_{n}\}$ with $i_{a}=\text{head}(p_{a})$:
$\mu_{S}:V_{{\boldsymbol{n}}}\rightarrow{\mathbb{Q}},\qquad\mu_{S}(F)=$
(60) $\underset{z_{n}=1}{\text{Res}}\dots\underset{z_{1}=1}{\text{Res}}\left(F(z_{1},\dots,z_{n})\frac{\prod_{1\leq a<b\leq n}(z_{b}-z_{a})^{\delta_{i_{a}i_{b}}-|\{\text{arrows }e=\overrightarrow{i_{a}i_{b}}|ep_{a}\leq p_{b}\}|}}{\prod_{a=1}^{n}(z_{a}-1)^{|\{1\leq\ell\leq m_{i_{a}}|o_{i_{a},\ell}\leq p_{a}\}|}}\right)$
Recall the lexicographic total order on acceptable collections, and let
[…]
Formula (56) is then an immediate consequence of the following
Claim 1: $\Theta_{{\boldsymbol{n}}}^{{\boldsymbol{m}}}=\bigcap_{S\text{ acceptable collection}}\text{Ker }\mu_{S}$
Claim 2: $\dim_{{\mathbb{Q}}}\left(V_{{\boldsymbol{n}},<S}\Big/V_{{\boldsymbol{n}},\leq S}\right)=1$ for any acceptable collection $S$.
Let us prove Claim 1: the inclusion $\subseteq$ is trivial, since the poles that arise in (52) are at least as bad as the poles that arise in (60). However, when we take in (60) the maximal acceptable collection $S=\{p_{1}<\dots<p_{n}\}$ with $\text{head}(p_{a})=i_{a}$, the poles in question are exactly the same as the poles that arise in (52). Therefore, the inclusion $\supseteq$ is an immediate consequence of the fact that
(61) $\mu_{S^{\prime}}(F)=0\quad\forall S^{\prime}\leq S\quad\Rightarrow\quad\mu_{S^{\prime}}(FG)=0\quad\forall S^{\prime}\leq S$
for all Laurent polynomials $F,G$ (in other words, $V_{{\boldsymbol{n}},\leq S}$ is an ideal). By induction on $S^{\prime}$, it suffices to show that $\mu_{S}(FG)=0$ for all Laurent polynomials $G$. By induction on $b\in\{1,\dots,n\}$, it suffices to consider (by Taylor expansion at $(1,\dots,1)$) either
  • $G=z_{b}-1$ if $p_{b}=o_{i_{b},\ell}$ for some $\ell$, or
  • $G=z_{b}-z_{a}$ if $p_{b}=ep_{a}$ for some $a<b$ and arrow $e=\overrightarrow{i_{a}i_{b}}$.
In either of these cases, we have
$\mu_{S}(FG)=\mu_{S^{\prime}}(FH)$
for some Laurent polynomial $H$, where $S^{\prime}$ is defined as the maximal acceptable collection with the same first b1b-1 paths as $S$, but whose $b$-th path is the maximal predecessor of $p_{b}$ in the set $\S(\{o_{i_{b},\ell}\}_{1\leq\ell\leq m_{i_{b}}}\cup\{ep_{a}\}_{a<b,e:i_{a}\rightarrow i_{b}})\backslash$ S. Since $S^{\prime}<S$, this completes the proof of (61).
In order to prove Claim 2, it is clear that the dimension in question is $\leq 1$, because any element of $V_{{\boldsymbol{n}},<S}/V_{{\boldsymbol{n}},\leq S}$ is uniquely determined by its image under $\mu_{S}$. To prove equality, it suffices to produce $F\in V_{{\boldsymbol{n}}}$ which is annihilated by all $\{\mu_{S^{\prime}}\}_{S^{\prime}<S}$ but is not annihilated by $\mu_{S}$. To this end, consider
$F_{S}=\text{Sym}\left[\frac{\prod_{a=1}^{n}(z_{a}-1)^{|\{1\leq\ell\leq m_{i_{a}}|o_{i_{a},\ell}<p_{a}\}|}\prod_{1\leq a<b\leq n}(z_{b}-z_{a})^{|\{e=\overrightarrow{i_{a}i_{b}}|ep_{a}<p_{b}\}|}}{\prod_{1\leq a<b\leq n,\ i_{a}=i_{b}}(z_{b}-z_{a})}\right]$
where “Sym” refers to averaging over all ways to permute pairs of variables $z_{a}$ and $z_{b}$ for which $i_{a}=i_{b}$. When evaluating $\mu_{S^{\prime}}(F_{S})$ for any $S^{\prime}=\{p_{1}^{\prime}<\dots<p_{n}^{\prime}\}$ using formula (60), the answer is equal to the sum of
$\underset{z_{n}=1}{\text{Res}}\dots\underset{z_{1}=1}{\text{Res}}\frac{\prod_{1\leq a<b\leq n}(z_{b}-z_{a})^{|\{\text{arrows }e|e\min(p_{\sigma(a)},p_{\sigma(b)})<\max(p_{\sigma(a)},p_{\sigma(b)})\}|-|\{e=\overrightarrow{i_{a}i_{b}}|ep_{a}^{\prime}\leq p_{b}^{\prime}\}|}}{\prod_{a=1}^{n}(z_{a}-1)^{|\{1\leq\ell\leq m_{i^{\prime}_{a}}|o_{i_{a}^{\prime},\ell}\leq p^{\prime}_{a}\}|-|\{\ell\in\{1,\dots,m_{i_{a}}\}|o_{i_{a}^{\prime},\ell}<p_{\sigma(a)}\}|}}$
as $\sigma\in S_{n}$ goes over all permutations such that $i_{\sigma(a)}=i^{\prime}_{a}$ for all $a\in\{1,\dots,n\}$ (we write $i_{a}^{\prime}=\text{head}(p_{a}^{\prime})$). It is clear that the residue above can be non-zero only if
  • $p_{1}^{\prime}>p_{\sigma(1)}$, or
  • $p_{1}^{\prime}=p_{\sigma(1)}$ and $p_{2}^{\prime}>p_{\sigma(2)}$, or
  • $p_{1}^{\prime}=p_{\sigma(1)}$, $p_{2}^{\prime}=p_{\sigma(2)}$, …, $p_{n-1}^{\prime}>p_{\sigma(n-1)}$, or
  • $p_{1}^{\prime}=p_{\sigma(1)}$, $p_{2}^{\prime}=p_{\sigma(2)}$, …, $p_{n-1}^{\prime}=p_{\sigma(n-1)}$, $p_{n}^{\prime}\geq p_{\sigma(n)}$.
If $S^{\prime}<S$, then the inequalities above are impossible, so $\mu_{S^{\prime}}(F_{S})=0$. Meanwhile if $S^{\prime}=S$, the inequalities above show that only the identity permutation $\sigma\in S_{n}$ produces a non-zero residue (whose value is $\pm 1$), and therefore $\mu_{S}(F_{S})\neq 0$. ∎
Proof.
of Lemma 2.13: If $\mathcal{T}$} is a rank $r$ vector bundle over a space $X$, we may define the projective space bundle ${\mathbb{P}}_{X}{\mathcal{T}}=\text{Proj}_{X}(\text{Sym}^{\bullet}({\mathcal{T}}))\rightarrow X$ which parameterizes one-dimensional quotients of $\mathcal{T}$}. It is well-known that
$K({\mathbb{P}}_{X}{\mathcal{T}})=K(X)[z^{\pm 1}]\Big/\wedge^{\bullet}\left(\frac{{\mathcal{T}}}{z}\right)$
where the variable $z$ has $K$-theoretic degree 0, and we write
(62) $\wedge^{\bullet}\left(\frac{{\mathcal{T}}}{z}\right)=\sum_{i=0}^{r}\frac{[\wedge^{i}{\mathcal{T}}]}{(-z)^{i}}\qquad\wedge^{\bullet}\left(\frac{z}{{\mathcal{T}}}\right)=\sum_{i=0}^{r}(-z)^{i}[\wedge^{i}{\mathcal{T}}^{\vee}]$
In other words, if we formally write ${\mathcal{T}}=\alpha_{1}+\dots+\alpha_{r}$, then we have
$K({\mathbb{P}}_{X}{\mathcal{T}})=K(X)[z^{\pm 1}]\Big/\left\{f(z)\Big|\sum_{k=1}^{r}\underset{z=\alpha_{k}}{\text{Res}}\left(\frac{z^{d}f(z)}{\wedge^{\bullet}({\mathcal{T}}/z)}\right)=0,\forall d\in{\mathbb{Z}}\right\}$
With this in mind, formula (59) reduces (via induction on $n$) to the claim that the forgetful map
$M^{\text{stab}}_{i_{1},\dots,i_{n},{\boldsymbol{m}}}\rightarrow M^{\text{stab}}_{i_{1},\dots,i_{n-1},{\boldsymbol{m}}}$
is the projectivization of a vector bundle with $K$-theory class equal to
[…]
where ${\mathcal{L}}_{1},\dots,{\mathcal{L}}_{n-1}$ are the line bundles on $M^{\text{stab}}_{i_{1},\dots,i_{n-1},{\boldsymbol{m}}}$ that parameterize the lines spanned by the vectors $v_{1},\dots,v_{n-1}$, respectively. The aforementioned claim is straightforward, and left to the reader. ∎
3. $q$-characters via shuffle algebras
We recall the shuffle algebra incarnation of quantum loop algebras $U_{q}(L{\mathfrak{g}})$ and of the simple modules $L({\boldsymbol{\psi}})$ in the Borel category $\mathcal{O}$}. The upshot will be a formula for the $q$-character $\chi_{q}(L^{\neq 0}({\boldsymbol{\psi}}))$ as the dimension of an explicit vector space of Laurent polynomials. In order to better compare the contents of the present Section with the more geometric Section 4, we work over the ground field ${\mathbb{K}}={\mathbb{Q}}(q)$, and consider only simple modules $U_{q}(L{\mathfrak{g}})\curvearrowright L({\boldsymbol{\psi}})$ with highest $\ell$-weight $\boldsymbol{\psi}$} of the form (77), i.e. the numerator and denominator of $\boldsymbol{\psi}$} completely factor into terms of the form $\{1-\gamma z^{-1}\}_{\gamma\in{\mathbb{K}}^{*}}$. This is equivalent to the standard treatment of the Borel category $\mathcal{O}$}, in which $\mathbb{K}$} is replaced by $\mathbb{C}$} and $q$ is replaced by a generic complex number.
3.1. Basic definitions
Fix a finite set $I$ and a generalized Cartan matrix
(63) $C=\left(c_{ij}=\frac{2d_{ij}}{d_{ii}}\in{\mathbb{Z}}\right)_{i,j\in I}$
where $d_{ii}=2d_{i}$ are positive integers with greatest common divisor 2, and $d_{ij}=d_{ji}$ are non-positive integers $\forall i\neq j$. This data corresponds to a symmetrizable Kac-Moody Lie algebra $\mathfrak{g}$}, which we will refer to as the “type” of all algebras considered hereafter. We interpret $I$ as a set of simple roots of $\mathfrak{g}$}, hence we interpret $\mathbb{Z}^{I}$} as the root lattice. We call $\mathfrak{g}$} finite type if the symmetric matrix $\left(d_{ij}\right)_{i,j\in I}$ is positive definite, since in this case $\mathfrak{g}$} is a finite-dimensional semisimple Lie algebra. Similarly, $\mathfrak{g}$} is called simply-laced if the generalized Cartan matrix (63) satisfies
(64) $d_{i}=1,\quad\ \forall i\in I$
Following the suggestion of David Hernandez, we call $\mathfrak{g}$} strongly symmetrizable if
(65) $d_{ij}\in\{0,-\max(d_{i},d_{j})\}$
for all $i\neq j$. All finite and affine type $\mathfrak{g}$} are strongly symmetrizable, except for $\widehat{{\mathfrak{sl}}}_{2}$.
3.2. The big shuffle algebra
We now review the trigonometric version ([14]) of the Feigin-Odesskii shuffle algebra ([15]) associated to the Kac-Moody Lie algebra $\mathfrak{g}$}. Consider the vector space of Laurent polynomials in arbitrarily many variables
(66) ${\mathcal{V}}=\bigoplus_{{\boldsymbol{n}}\in{\mathbb{N}^{I}}}{\mathcal{V}}_{{\boldsymbol{n}}},\quad\text{where}\quad{\mathcal{V}}_{(n_{i}\geq 0)_{i\in I}}={\mathbb{K}}[z_{i1}^{\pm 1},\dots,z_{in_{i}}^{\pm 1}]^{\text{sym}}_{i\in I}$
Above, “sym” refers to color-symmetric Laurent polynomials, meaning that they are symmetric in the variables $z_{i1},\dots,z_{in_{i}}$ for each $i\in I$ separately (the terminology is inspired by the fact that $i\in I$ is called the color of the variable $z_{ia}$). We make the vector space $\mathcal{V}$} into a $\mathbb{K}$}-algebra via the following shuffle product:
(67) $F(z_{i1},\dots,z_{in_{i}})_{i\in I}*F^{\prime}(z_{i1},\dots,z_{in^{\prime}_{i}})_{i\in I}=\frac{1}{{\boldsymbol{n}}!{\boldsymbol{n}}^{\prime}!}\,\cdot$
$\textrm{Sym}\left[F(z_{i1},\dots,z_{in_{i}})F^{\prime}(z_{i,n_{i}+1},\dots,z_{i,n_{i}+n^{\prime}_{i}})\prod_{i,j\in I}\mathop{\prod_{1\leq a\leq n_{i}}}_{n_{j}<b\leq n_{j}+n_{j}^{\prime}}\zeta_{ji}\left(\frac{z_{jb}}{z_{ia}}\right)\right]$
where we fix a total order << on $I$ and define  [^4] Note that the function $\zeta_{ij}$ below is equal to $1-x^{\delta_{i<j}-\delta_{i>j}}$ times the same-named function studied in [38]. This means that the algebra $\mathcal{V}$} of (66) is isomorphic to the same-named algebra in loc. cit., with the isomorphism given by multiplication with $\prod_{i<j}\prod_{a=1}^{n_{i}}\prod_{b=1}^{n_{j}}\left(1-\frac{z_{ia}}{z_{jb}}\right)$.
(68) $\zeta_{ij}(x)=\begin{cases}q^{-d_{ij}}-x&\text{if }i<j\\ \frac{q^{-d_{ij}}-x}{1-x}&\text{if }i=j\\ 1-q^{-d_{ij}}x^{-1}&\text{if }i>j\end{cases}$
The word “Sym” in (67) denotes symmetrization with respect to the
$({\boldsymbol{n}}+{\boldsymbol{n}}^{\prime})!:=\prod_{i\in I}(n_{i}+n^{\prime}_{i})!$
permutations of the variables $\{z_{i1},\dots,z_{i,n_{i}+n^{\prime}_{i}}\}$ for each $i$ independently. The reason for formula (67) is to ensure that there exists an algebra homomorphism
(69) $\Upsilon:U_{q}^{-}(L{\mathfrak{g}})\rightarrow{\mathcal{V}}$
whose domain is the negative half (in Drinfeld’s new realization, [11]) of the quantum loop algebra $U_{q}(L{\mathfrak{g}})$ of $\mathfrak{g}$}. We refer to [38] for an overview of the generators-and-relations presentation of $U_{q}(L{\mathfrak{g}})$, and note that (by definition) it involves sufficiently many relations so that $\Upsilon$ is injective. The corresponding set of relations is so far only known explicitly for finite type ([2, 11]), more generally for strongly symmetrizable ([40]), and for simply-laced ([36]) symmetrizable Kac-Moody Lie algebras $\mathfrak{g}$}. See [37] for a discussion.
3.3. The (small) shuffle algebra
We will refer to
(70) ${\mathcal{S}}=\text{Im }\Upsilon$
as simply the shuffle algebra, i.e. the subalgebra of $\mathcal{V}$} generated by $\{z_{i1}^{d}\in{\mathcal{V}}_{\boldsymbol{\varsigma}^{i}}\}_{i\in I,d\in{\mathbb{Z}}}$. Note that $\mathcal{S}$} was denoted by $\mathcal{S}$$^{-}$ in [38, 39], in order to differentiate it from its opposite algebra ${\mathcal{S}}^{\text{op}}={\mathcal{S}}^{+}$; this distinction will not play any role in the present paper. Explicitly, $\mathcal{S}$} is the $\mathbb{K}$}-linear span of Laurent polynomials of the form
(71) $\text{Sym}\left[z_{1}^{d_{1}}\dots z_{n}^{d_{n}}\prod_{1\leq a<b\leq n}\zeta_{i_{b}i_{a}}\left(\frac{z_{b}}{z_{a}}\right)\right]$
as $i_{1},\dots,i_{n}$ run over orderings of $\boldsymbol{n}$}, and $d_{1},\dots,d_{n}$ run over $\mathbb{Z}$} (in the above expression, we identify each variable $z_{a}$ with some variable of the form $z_{i_{a}\bullet_{a}}$ in the notation of (66); the values of the numbers $\bullet_{a}$ do not matter due to the presence of the symmetrization, as long as we have $\bullet_{a}\neq\bullet_{b}$ whenever $a\neq b$, $i_{a}=i_{b}$).
Example 3.4.
For $\mathfrak{g}$} of finite type, we have the following explicit description
[…]
As $i\neq j$ range over $I$, the following are called Feigin-Odesskii wheel conditions:
(73) $F(\dots,z_{ia},\dots,z_{jb},\dots)\Big|_{(z_{i1},z_{i2},\dots,z_{i,1-c_{ij}})\mapsto(z_{j1}q^{d_{ij}},z_{j1}q^{d_{ij}+d_{ii}},\dots,z_{j1}q^{-d_{ij}})}=0$
The inclusion $\subseteq$ of (72) was established in [14] based on the seminal work [15], while the inclusion $\supseteq$ of (72) was proved in [41]. More generally, the formulas above hold as stated for strongly symmetrizable $\mathfrak{g}$}, as shown in [40, Section 3.25].
Example 3.5.
For any simply-laced Kac-Moody $\mathfrak{g}$}, it was shown in [36] that $\mathcal{S}$} is the set of color-symmetric Laurent polynomials $F$ that satisfy the following condition: for any $i\neq j$ in $I$ and any $t-s,t^{\prime}-s^{\prime}\in 2{\mathbb{N}}$ such that $s-t^{\prime}=s^{\prime}-t\equiv d_{ij}$ mod 2,
(74) $(x-y)^{\frac{t-s^{\prime}+d_{ij}}{2}+1}$
divides the specialization of $F$ at
(75) $z_{i1}=xq^{s},\ \quad z_{i2}=xq^{s+2},\quad\ \dots\ \quad z_{i,\frac{t-s}{2}}=xq^{t-2},\ \ \quad z_{i,\frac{t-s}{2}+1}=xq^{t}$
(76) $z_{j1}=yq^{s^{\prime}},\quad z_{j2}=yq^{s^{\prime}+2},\quad\dots\quad z_{j,\frac{t^{\prime}-s^{\prime}}{2}}=yq^{t^{\prime}-2},\quad z_{j,\frac{t^{\prime}-s^{\prime}}{2}+1}=yq^{t^{\prime}}$
3.6. Simple modules and $q$-characters
The isomorphism between $U_{q}(L{\mathfrak{g}})$ and the double of the shuffle algebra $\mathcal{S}$} allowed us to give the following explicit construction of an important class of simple modules ([38]). For any highest $\ell$-weight
(77) ${\boldsymbol{\psi}}=\left(\psi_{i}(z)=\text{constant}_{i}\cdot\frac{\prod_{\gamma\in{\mathbb{K}}^{*}}\left(1-\frac{\gamma}{z}\right)^{l_{i,\gamma}}}{\prod_{\gamma\in{\mathbb{K}}^{*}}\left(1-\frac{\gamma}{z}\right)^{m_{i,\gamma}}}\right)_{i\in I}$
with various $l_{i,\gamma},m_{i,\gamma}\geq 0$ (almost all of which are actually equal to 0) one considers the corresponding simple module $L({\boldsymbol{\psi}})$ of the Borel half of $U_{q}(L{\mathfrak{g}})$ (defined in [25] in the finite type case and in [38] in the arbitrary symmetrizable Kac-Moody case). It was shown in [39, Proposition 4.7] that there is a isomorphism of vector spaces
(78) $L({\boldsymbol{\psi}})\cong L^{\textbf{ord }{\boldsymbol{\psi}}}\otimes L^{\neq 0}({\boldsymbol{\psi}})$
The first tensor factor in the RHS above was calculated in [39] for finite type $\mathfrak{g}$} (proving a conjecture of [31, 49]) and we will not be concerned with it. Instead, we will focus our attention on the second tensor factor in the RHS of (78): this factor was realized in [39, Definition 4.3, Remark 4.5] as
(79) $L^{\neq 0}({\boldsymbol{\psi}})=\bigoplus_{{\boldsymbol{n}}\in{\mathbb{N}^{I}}}{\mathcal{S}}_{{\boldsymbol{n}}}\Big/\left({\mathcal{S}}_{{\boldsymbol{n}}}\cap\Theta({\boldsymbol{\psi}})_{{\boldsymbol{n}}}\right)$
where $\Theta({\boldsymbol{\psi}})_{{\boldsymbol{n}}}$ denotes the set of $F(z_{i1},\dots,z_{in_{i}})_{i\in I}\in{\mathcal{V}}_{{\boldsymbol{n}}}$ such that  [^5] Below, the symbol $\underset{z\in{\mathbb{K}}^{*}}{\text{Res}}$” should be interpreted as “the sum over all residues in $\mathbb{K}$$^{*}$”.
(80) $\underset{z_{n}\in{\mathbb{K}}^{*}}{\text{Res}}\dots\underset{z_{1}\in{\mathbb{K}}^{*}}{\text{Res}}\left(\frac{z_{1}^{d_{1}}\dots z_{n}^{d_{n}}F(z_{1},\dots,z_{n})}{\prod_{1\leq a<b\leq n}\zeta_{i_{b}i_{a}}\left(\frac{z_{b}}{z_{a}}\right)}\prod_{a=1}^{n}\frac{\psi_{i_{a}}(z_{a})}{z_{a}}\right)=0$
for all orderings $i_{1},\dots,i_{n}$ of $\boldsymbol{n}$} and all $d_{1},\dots,d_{n}\in{\mathbb{Z}}$. Given such an ordering, the notation $F(z_{1},\dots,z_{n})$ in the RHS of (80) refers to the fact that each symbol $z_{a}$ is matched with a variable of the form $z_{i_{a}\bullet_{a}}$ of $F$, where the choice of $\bullet_{a}$ does not matter due to the color-symmetry of $F$ (as long as $\bullet_{a}\neq\bullet_{b}$ whenever $a\neq b,i_{a}=i_{b}$).
Since both ${\mathcal{S}}_{{\boldsymbol{n}}}$ and $\Theta({\boldsymbol{\psi}})_{{\boldsymbol{n}}}$ are preserved by multiplication with arbitrary elements of the ring ${\mathcal{P}}_{\boldsymbol{n}}^{\pm}$ of color-symmetric Laurent polynomials, we can decompose the right-hand side of (79) in terms of its fibers over the maximal spectrum
$\text{Spec}({\mathcal{P}}^{\pm}_{{\boldsymbol{n}}})\supseteq({\mathbb{K}}^{*})^{{\boldsymbol{n}}}:=\prod_{i\in I}({\mathbb{K}}^{*})^{n_{i}}/S_{n_{i}}$
Explicitly, as $\boldsymbol{x}=(x_{ia})_{i\in I,a\in\{1,\dots,n_{i}\}}$ runs over $({\mathbb{K}}^{*})^{{\boldsymbol{n}}}$, we have
(81) $L^{\neq 0}({\boldsymbol{\psi}})=\mathop{\bigoplus_{{\boldsymbol{n}}\in{\mathbb{N}^{I}}}}_{\boldsymbol{x}\in({\mathbb{K}}^{*})^{{\boldsymbol{n}}}}\left({\mathcal{S}}_{{\boldsymbol{n}}}\Big/\left({\mathcal{S}}_{{\boldsymbol{n}}}\cap\Theta({\boldsymbol{\psi}})_{{\boldsymbol{n}}}\right)\right)_{\boldsymbol{x}}=\mathop{\bigoplus_{{\boldsymbol{n}}\in{\mathbb{N}^{I}}}}_{\boldsymbol{x}\in({\mathbb{K}}^{*})^{{\boldsymbol{n}}}}{\mathcal{S}}_{{\boldsymbol{n}}}\Big/\left({\mathcal{S}}_{{\boldsymbol{n}}}\cap\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}}\right)$
where $\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}}$ denotes the set of $F\in{\mathcal{V}}_{{\boldsymbol{n}}}$ such that
(82) $\underset{z_{n}=x_{n}}{\text{Res}}\dots\underset{z_{1}=x_{1}}{\text{Res}}\left(\frac{z_{1}^{d_{1}}\dots z_{n}^{d_{n}}F(z_{1},\dots,z_{n})}{\prod_{1\leq a<b\leq n}\zeta_{i_{b}i_{a}}\left(\frac{z_{b}}{z_{a}}\right)}\prod_{a=1}^{n}\frac{\psi_{i_{a}}(z_{a})}{z_{a}}\right)=0$
for any ordering $x_{1},\dots,x_{n}$ of (the entries of the tuple) $\boldsymbol{x}$ underlain by any ordering $i_{1},\dots,i_{n}$ of $\boldsymbol{n}$} (in other words, each $x_{a}$ is a placeholder for one of the coordinates $x_{i_{a}\bullet_{a}}$ of $\boldsymbol{x}$, where the numbers $\bullet_{a}$ are arbitrary as long as $\bullet_{a}\neq\bullet_{b}$ if $a\neq b$, $i_{a}=i_{b}$).
3.7. A model for shuffle algebras
As explained in Subsection 1.2, we are interested in finding a geometric incarnation for the non-negative integers
(83) $\dim_{{\mathbb{K}}}\left({\mathcal{S}}_{{\boldsymbol{n}}}\Big/\left({\mathcal{S}}_{{\boldsymbol{n}}}\cap\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}}\right)\right)$
for any (henceforth fixed) ${\boldsymbol{n}}\in{\mathbb{N}^{I}}$ and $\boldsymbol{x}\in({\mathbb{K}}^{*})^{{\boldsymbol{n}}}$. We have the following inclusion
(84) ${\mathcal{S}}_{{\boldsymbol{n}}}\Big/\left({\mathcal{S}}_{{\boldsymbol{n}}}\cap\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}}\right)\hookrightarrow{\mathcal{V}}_{{\boldsymbol{n}}}\Big/\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}}$
and we will proceed to give a purely combinatorial description of the vector space in the right-hand side of (84). Define
(85) $V_{\boldsymbol{x}}={\mathbb{Q}}[z_{i1}^{\pm 1},\dots,z_{in_{i}}^{\pm 1}]_{i\in I}^{\text{sym}_{\boldsymbol{x}}}$
where $\textrm{sym}_{\boldsymbol{x}}$ refers to the partial symmetrization that only permutes those $z_{ia}$ and $z_{ib}$ for which $x_{ia}=x_{ib}$. For any tuple of integers ${\boldsymbol{m}}=(m_{ia})_{i\in I,1\leq a\leq n_{i}}$, let
$\Theta_{\boldsymbol{x}}^{{\boldsymbol{m}}}\subseteq V_{\boldsymbol{x}}$
be the intersection over all orderings $x_{1},\dots,x_{n}$ of the entries of $\boldsymbol{x}$ (with underlying ordering $i_{1},\dots,i_{n}$ of $\boldsymbol{n}$}) of the sets of Laurent polynomials $F\in V_{\boldsymbol{x}}$ such that
(86) $\underset{z_{n}=1}{\text{Res}}\dots\underset{z_{1}=1}{\text{Res}}\left(\frac{z_{1}^{d_{1}}\dots z_{n}^{d_{n}}F(z_{1},\dots,z_{n})}{\prod_{a=1}^{n}(z_{a}-1)^{m_{a}}}\right.\\ \left.\prod_{1\leq a<b\leq n}(z_{b}-z_{a})^{\delta_{(i_{a},x_{a})=(i_{b},x_{b})}-\delta_{x_{a}=x_{b}q^{d_{i_{a}i_{b}}}}}\right)=0$
for all $d_{1},\dots,d_{n}\in{\mathbb{Z}}$, where the numbers $m_{1},\dots,m_{n}$ represent an ordering of the entries of $\boldsymbol{m}$} in the same way as $x_{1},\dots,x_{n}$ represent an ordering of the entries of $\boldsymbol{x}$.
Remark 3.8.
If $m_{ia}\geq 0,\forall i,a$, then $\Theta_{\boldsymbol{x}}^{{\boldsymbol{m}}}$ coincides with (55) for the quiver with
  • vertex set $I_{\boldsymbol{x}}=\left\{(i,a)\Big|i\in I,1\leq a\leq n_{i}\right\}\Big/\Big((i,a)\sim(i,b)\text{ if }x_{ia}=x_{ib}\Big)$;
  • arrow set $(i,a)\rightarrow(j,b)$ whenever $x_{ia}=x_{jb}q^{d_{ij}}$;
  • dimension vector $(i,a)\leadsto$ number of times $x_{ia}$ appears in the multiset $\{x_{i1},\dots,x_{in_{i}}\}$
  • framing dimension $m_{ia}$ at any vertex $(i,a)$.
Proposition 3.9.
We have an isomorphism of finite-dimensional vector spaces
(87) ${\mathcal{V}}_{{\boldsymbol{n}}}\Big/\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}}\cong\left(V_{\boldsymbol{x}}\Big/\Theta^{{\boldsymbol{m}}}_{\boldsymbol{x}}\right)\otimes_{{\mathbb{Q}}}{\mathbb{K}}$
where the tuple $\boldsymbol{m}$} is related to $\boldsymbol{\psi}$} by the formula $\psi_{i}(z)=(z-x_{ia})^{-m_{ia}}$ times a rational function which is regular and non-zero at $z=x_{ia}$.
Proof.
There is a natural map
(88) ${\mathcal{V}}_{{\boldsymbol{n}}}\hookrightarrow V_{\boldsymbol{x}}\otimes_{{\mathbb{Q}}}{\mathbb{K}}$
defined by rescaling the variables of Laurent polynomials according to
$z_{ia}\leadsto z_{ia}x_{ia}$
and then regarding color-symmetric Laurent polynomials as being simply partially symmetric with respect to those $z_{ia}$ and $z_{ib}$ such that $x_{ia}=x_{ib}$. To show that the assignment (88) sends $\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}}$ to $\Theta_{\boldsymbol{x}}^{{\boldsymbol{m}}}$, we will argue as follows: a Laurent polynomial $F$ lies inside $\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}}$ if and only if finitely many linear combinations of its derivatives vanish at $\boldsymbol{x}$. This vanishing property is preserved upon multiplying $F$ by any rational function which is regular and non-zero at $\boldsymbol{x}$. However, this is precisely the relation between the integrands of (82) and (86): the latter expression involves the linear factors of the former expression which vanish at the point $\boldsymbol{x}$ and nothing more. Thus, $F$ lies in $\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}}$ if and only if its image under (88) lies in $\Theta_{\boldsymbol{x}}^{{\boldsymbol{m}}}$.
The previous argument precisely shows that (88) gives rise to an injective map
(89) ${\mathcal{V}}_{{\boldsymbol{n}}}\Big/\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}}\hookrightarrow\left(V_{\boldsymbol{x}}\Big/\Theta^{{\boldsymbol{m}}}_{\boldsymbol{x}}\right)\otimes_{{\mathbb{Q}}}{\mathbb{K}}$
We will show that (89) is also surjective by the following general argument. Consider any parabolic subgroup (in the case at hand, $P$ will be the stabilizer of $\boldsymbol{x}$)
$P=S_{n_{1}}\times\dots\times S_{n_{k}}\hookrightarrow S_{n}$
where $n=n_{1}+\dots+n_{k}$, and any point
$x=(x_{1},\dots,x_{n})\in({\mathbb{C}}^{*})^{n}$
such that $x_{i}=x_{j}$ if and only if $(ij)\in P$. Then for any $P$-invariant Laurent polynomial $f(z_{1},\dots,z_{n})$, we claim that there exists a $S_{n}$-invariant Laurent polynomial $g$ which has the same $N$-th order (for arbitrarily large $N$) derivatives at $x$ as $f$; it is clear that the underlined claim precisely establishes the surjectivity of (89). To prove the underlined claim, it suffices to take
$g=\sum_{\sigma\in S_{n}/S_{n_{1}}\times\dots\times S_{n_{k}}}\sigma(f\cdot r)$
where $r(z_{1},\dots,z_{n})$ is a $P$-invariant Laurent polynomial such that
$r\in\Big(1+\left((z_{1}-x_{1})^{N},\dots,(z_{n}-x_{n})^{N}\right)\Big)\bigcap_{(ij)\notin P}\Big((z_{i}-x_{j})^{N}\Big)$
The existence of such a Laurent polynomial $r$ is due to the fact that the polynomial
$p(z_{1},\dots,z_{n})=\prod_{(ij)\notin P}(z_{i}-x_{j})^{N}$
does not vanish at $(z_{1},\dots,z_{n})=(x_{1},\dots,x_{n})$, and thus the principal ideal $(p)$ together with $((z_{1}-x_{1})^{N},\dots,(z_{n}-x_{n})^{N})$ generate the unit ideal. ∎
4. Relation to critical $K$-theory
In Subsections 1.2 and 3.6, we saw that the coefficients of $q$-characters of simple modules $L({\boldsymbol{\psi}})$ can be obtained from the dimensions of the vector spaces
(90) ${\mathcal{S}}_{{\boldsymbol{n}}}\Big/\left({\mathcal{S}}_{{\boldsymbol{n}}}\cap\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}}\right)$
as $\boldsymbol{n}$} ranges over $\mathbb{N}^{I}$} and $\boldsymbol{x}$ ranges over $({\mathbb{K}}^{*})^{{\boldsymbol{n}}}$. We will now relate the vector spaces (90) to critical $K$-theory of quiver varieties, which we will argue provides a suitable generalization of Theorem 1.4 beyond reachable simple modules. We assume that $\mathfrak{g}$} is strongly symmetrizable throughout the present Section; recall that this includes all finite and affine type symmetrizable Kac-Moody Lie algebras except for $\widehat{{\mathfrak{sl}}}_{2}$.
4.1. Quiver with potential
Let us recall from Subsection 1.7 the quiver $Q$ with
  • vertex set $I$,
  • arrows $j\xleftarrow{{}_{j}\square_{i}}i$ for all $i,j\in I$ (including $i=j$),
  • potential (i.e. a linear combination of cycles in the path algebra of $Q$)
    (91) $W=\sum_{i\neq j}{{}_{j}\square_{i}}\underbrace{{{}_{i}\square_{i}}\dots{}_{i}\square_{i}}_{-c_{ij}\text{ factors}}{{}_{i}\square_{j}}$
    where $c_{ij}=\frac{2d_{ij}}{d_{ii}}\in{\mathbb{Z}}_{\leq 0}$ are the off-diagonal coefficients of the Cartan matrix.
In [47, 48], the authors considered the critical $K$-theory of the above quiver with potential (following [1, 12, 28], see also [43]). We refer to the aforementioned works for an introduction to critical $K$-theory, and restrict ourselves to simply listing the following key facts that will be used in the present paper.
  • For any ${\boldsymbol{n}}\in{\mathbb{N}^{I}}$, recall the stack $M_{{\boldsymbol{n}}}$ of $\boldsymbol{n}$}-dimensional representations of $Q$. There is an action ${\mathbb{C}}^{*}\curvearrowright M_{{\boldsymbol{n}}}$ which assigns weight $q^{d_{ij}}$ to any arrow $_{j}\square_{i}$.
  • Evaluating the trace of the potential $W$ in quiver representations gives a function
    (92) $\text{Tr}(W):M_{{\boldsymbol{n}}}\rightarrow{\mathbb{C}}$
[…]
    (93) $K^{{\mathbb{C}}^{*}}(Q,W)$
[…]
  • The critical $K$-theory (93) is a module over the equivariant $K$-theory of the stack
    $K^{{\mathbb{C}}^{*}}(M_{{\boldsymbol{n}}})={\mathbb{K}}[z_{i1}^{\pm 1},\dots,z_{in_{i}}^{\pm 1}]_{i\in I}^{\text{sym}}={\mathcal{V}}_{{\boldsymbol{n}}}$
  • There is a ${\mathcal{V}}_{{\boldsymbol{n}}}$-module homomorphism
    (94) $K^{{\mathbb{C}}^{*}}(Q,W)_{{\boldsymbol{n}}}\xrightarrow{\iota_{{\boldsymbol{n}}}}{\mathcal{V}}_{{\boldsymbol{n}}}$
    ([42, Proposition 3.6]) whose image is contained in the image of the push-forward
    (95) $K^{{\mathbb{C}}^{*}}(Z_{{\boldsymbol{n}}})\xrightarrow{\alpha}K^{{\mathbb{C}}^{*}}(M_{{\boldsymbol{n}}})$
    with $Z_{{\boldsymbol{n}}}$ denoting the zero locus of the function $\text{Tr}(W):M_{{\boldsymbol{n}}}\rightarrow{\mathbb{C}}$.
  • There is a convolution algebra structure ([47, 48], see also [43]) on
    (96) $K^{{\mathbb{C}}^{*}}(Q,W)=\bigoplus_{{\boldsymbol{n}}\in{\mathbb{N}^{I}}}K^{{\mathbb{C}}^{*}}(Q,W)_{{\boldsymbol{n}}}$
    that is referred to as the critical convolution algebra. The map
    (97) $K^{{\mathbb{C}}^{*}}(Q,W)\xrightarrow{\iota}{\mathcal{V}}$
    obtained by putting together all the maps (94) is an algebra homomorphism, where $\mathcal{V}$} is made into an algebra with respect to the shuffle product (67).
The above notions are particularly simple when ${\boldsymbol{n}}=\boldsymbol{\varsigma}^{i}$ for some $i\in I$. In this case, the moduli stack of quiver representations is $\mathbb{C}$$\times\cdot$$\mathbb{C}$$^{*}$, and so its $K$-theory is
$K^{{\mathbb{C}}^{*}}(Q,W)_{\boldsymbol{\varsigma}^{i}}=K^{{\mathbb{C}}^{*}}(M_{\boldsymbol{\varsigma}^{i}})\cong{\mathbb{K}}[z_{i1}^{\pm 1}]={\mathcal{V}}_{\boldsymbol{\varsigma}^{i}}$
In particular, since the $\mathbb{K}$}-algebra $\mathcal{S}$} is generated by $\{z_{i1}^{d}\}_{i\in I,d\in{\mathbb{Z}}}$, we have
(98) $\text{Im }\iota\supseteq{\mathcal{S}}$
Proposition 4.2.
If $\mathfrak{g}$} is strongly symmetrizable, then $\emph{Im }\iota={\mathcal{S}}$.
Proof.
We follow the idea of Yu Zhao from [50]. Recall the explicit description of $\mathcal{S}$} from Example 3.4. It suffices to prove that $\text{Im }\alpha\subseteq{\mathcal{S}}_{{\boldsymbol{n}}}$ with $\alpha$ being the map (95). Because of the localization exact sequence
[…]
we must prove that $\text{Ker }\beta\subseteq{\mathcal{S}}_{{\boldsymbol{n}}}$. To this end, fix $i\neq j$ in $I$ and consider the locally closed substack $U\subset M_{{\boldsymbol{n}}}$ that parameterizes quiver representations of the form
[…]
$U\cong({\mathbb{C}}^{*})^{2-c_{ij}}\Big/({\mathbb{C}}^{*})^{2-c_{ij}}\cong{\mathbb{C}}^{*}\times\left(\cdot/{\mathbb{C}}^{*}\right)$
where the left-most $({\mathbb{C}}^{*})^{2-c_{ij}}$ parameterizes the arrow maps in (99) and the right-most $({\mathbb{C}}^{*})^{2-c_{ij}}$ acts on it by conjugation. Restricting a Laurent polynomial in ${\mathcal{V}}_{{\boldsymbol{n}}}\cong K^{{\mathbb{C}}^{*}}(M_{{\boldsymbol{n}}})$ to $U$ has the effect of performing the specialization (73).
4.3. Graded quiver with potential
Let us now consider the graded quiver $Q^{\text{gr}}$ of Subsection 1.3, which has
  • vertex set $I\times{\mathbb{Z}}$,
  • arrows $(j,t)\xleftarrow{{}^{t}_{j}\square^{s}_{i}}(i,s)$ for all $i,j\in I$ and $s,t\in{\mathbb{Z}}$ with $t=s-d_{ij}$;
  • potential
    $W^{\text{gr}}=\sum_{i\neq j}\sum_{s\in{\mathbb{Z}}}\left({}^{s}_{j}\square_{i}^{s+d_{ij}}\right)\left({}^{s+d_{ij}}_{i}\square_{i}^{s+d_{ij}+d_{ii}}\right)\dots\left({}^{s-d_{ij}-d_{ii}}_{i}\square_{i}^{s-d_{ij}}\right)\left({}^{s-d_{ij}}_{i}\square_{j}^{s}\right)$
The connection between the graded quiver above and the construction of Subsection 4.1 is given via the fixed points of the action ${\mathbb{C}}^{*}\curvearrowright M_{{\boldsymbol{n}}}$. More specifically, we have
(100) $M_{{\boldsymbol{n}}}^{{\mathbb{C}}^{*}}\supset\bigsqcup_{\boldsymbol{x}\in q^{{\mathbb{Z}}^{{\boldsymbol{n}}}}}M_{\boldsymbol{x}}$
where in the RHS, $M_{\boldsymbol{x}}$ denotes the stack of $Q^{\text{gr}}$-representations which have dimension at $(i,s)$ equal to the number of times $q^{s}$ appears among the coordinates of $\boldsymbol{x}=(x_{ia})_{i\in I,1\leq a\leq n_{i}}$. The map (88) can be interpreted as the restriction from the $K$-theory of $M_{{\boldsymbol{n}}}$ to that of $M_{\boldsymbol{x}}$ (note that $K(M_{\boldsymbol{x}})=V_{\boldsymbol{x}}$). We have the following non-equivariant analogue of the objects in (94)
(101) $K(Q^{\text{gr}},W^{\text{gr}})_{\boldsymbol{x}}\xrightarrow{\iota^{\text{gr}}_{\boldsymbol{x}}}V_{\boldsymbol{x}}$
4.4. Framed critical $K$-theory
Let us henceforth consider a highest $\ell$-weight $\boldsymbol{\psi}$} of the form
${\boldsymbol{\psi}}=\left(\psi_{i}(z)=\text{constant}_{i}\cdot\frac{\psi_{i}^{\text{num}}(z)}{\psi_{i}^{\text{den}}(z)}\right)_{i\in I}$
as in (7). In other words, we have that ${\boldsymbol{\psi}}^{\text{num}}=(\psi_{i}^{\text{num}}(z))_{i\in I}$, ${\boldsymbol{\psi}}^{\text{den}}=(\psi_{i}^{\text{den}}(z))_{i\in I}$ are Laurent polynomial $\ell$-weights of the form
$\psi_{i}^{\text{num}}(z)=\prod_{s\in{\mathbb{Z}}}\left(1-\frac{q^{s}}{z}\right)^{l_{i,s}}$
$\psi_{i}^{\text{den}}(z)=\prod_{s\in{\mathbb{Z}}}\left(1-\frac{q^{s}}{z}\right)^{m_{i,s}}$
for various $l_{i,s},m_{i,s}\in{\mathbb{N}}$, almost all of which are 0. We will write ${\mathbf{l}}=(l_{i,s})_{i\in I,s\in{\mathbb{Z}}}$ and ${\boldsymbol{m}}=(m_{i,s})_{i\in I,s\in{\mathbb{Z}}}$. Consider the quivers from the previous Subsections, but framed as follows:
  • we endow $Q$ with framing $\oplus_{s\in{\mathbb{Z}}}{\mathbb{C}}^{m_{i,s}}$ at every vertex $i\in I$, and declare the $s$-th direct summand to be equivariant with respect to the character $q^{s}$ of $\mathbb{C}$$^{*}$.
  • we endow $Q^{\text{gr}}$ with framing ${\mathbb{C}}^{m_{i,s}}$ at every vertex $(i,s)\in I\times{\mathbb{Z}}$.
The choices above allow to define the stacks of framed quiver representations $M_{{\boldsymbol{n}},{\boldsymbol{m}}}$ and $M_{\boldsymbol{x},{\boldsymbol{m}}}$, which are total spaces of affine bundles over the stacks $M_{{\boldsymbol{n}}}$ and $M_{\boldsymbol{x}}$, respectively. As such, the corresponding framed critical $K$-theory groups are identified with the unframed ones, i.e. in the following equations
(102) $K^{{\mathbb{C}}^{*}}(Q,W)_{{\boldsymbol{n}},{\boldsymbol{m}}}\xrightarrow{\iota_{{\boldsymbol{n}},{\boldsymbol{m}}}}{\mathcal{V}}_{{\boldsymbol{n}}}$
(103) $K(Q^{\text{gr}},W^{\text{gr}})_{\boldsymbol{x},{\boldsymbol{m}}}\xrightarrow{\iota^{\text{gr}}_{\boldsymbol{x},{\boldsymbol{m}}}}V_{\boldsymbol{x}}$
both the LHS and the map are identified with those of (94) and (101), respectively.
The Laurent polynomial $\ell$-weight ${\boldsymbol{\psi}}^{\text{num}}$ gives rise to the following element
(104) $\Delta_{{\mathbf{l}}}=\prod_{i\in I}\prod_{s\in{\mathbb{Z}}}\wedge^{\bullet}\left(\frac{q^{s}}{{\mathcal{T}}_{i}}\right)^{l_{i,s}}$
in the $K$-theory of $M_{{\boldsymbol{n}},{\boldsymbol{m}}}$ (a.k.a. ${\mathcal{V}}_{{\boldsymbol{n}}}$), where ${\mathcal{T}}_{i}$ denotes the tautological vector bundle corresponding to the vertex $i\in I$. Similarly, we have the element
(105) $\Delta^{\text{gr}}_{{\mathbf{l}}}=\prod_{i\in I}\prod_{s\in{\mathbb{Z}}}\wedge^{\bullet}\left(\frac{1}{T_{i,s}}\right)^{l_{i,s}}$
in the $K$-theory of $M_{\boldsymbol{x},{\boldsymbol{m}}}$ (a.k.a. $V_{\boldsymbol{x}}$), where $T_{i,s}$ denotes the tautological vector bundle corresponding to the vertex $(i,s)\in I\times{\mathbb{Z}}$. The relation between $\Delta_{{\mathbf{l}}}$ and $\Delta_{{\mathbf{l}}}^{\text{gr}}$ is that the latter is the product of those linear factors of the former which do not involve any $q$’s, upon invoking the restriction map
[…]
Lemma 4.5.
Under the maps (102) and (103), the operations
$K^{{\mathbb{C}}^{*}}(Q,W)_{{\boldsymbol{n}},{\boldsymbol{m}}}\xrightarrow{\cdot\Delta_{{\mathbf{l}}}}K^{{\mathbb{C}}^{*}}(Q,W)_{{\boldsymbol{n}},{\boldsymbol{m}}}$
$K(Q^{\emph{gr}},W^{\emph{gr}})_{\boldsymbol{x},{\boldsymbol{m}}}\xrightarrow{\cdot\Delta_{{\mathbf{l}}}^{\emph{gr}}}K(Q^{\emph{gr}},W^{\emph{gr}})_{\boldsymbol{x},{\boldsymbol{m}}}$
correspond to multiplication by elements which we abusively denote by
(106) $\Delta_{{\mathbf{l}}}=\prod_{i\in I}\prod_{a=1}^{n_{i}}\prod_{s\in{\mathbb{Z}}}\left(1-\frac{q^{s}}{z_{ia}}\right)^{l_{i,s}}\in{\mathcal{V}}_{{\boldsymbol{n}}}$
(107) $\Delta^{\emph{gr}}_{{\mathbf{l}}}=\prod_{i\in I}\prod_{a=1}^{n_{i}}\left(1-\frac{1}{z_{ia}}\right)^{l_{i,\log_{q}(x_{ia})}}\in V_{\boldsymbol{x}}$
respectively. The relation between $\Delta_{{\mathbf{l}}}$ and $\Delta_{{\mathbf{l}}}^{\emph{gr}}$ is that the latter consists of all the linear factors of the former which vanish at $\{z_{ia}=x_{ia}\}_{i\in I,1\leq a\leq n_{i}}$  [^6] The apparent mismatch in notation between $\Delta_{{\mathbf{l}}}$ and $\Delta_{{\mathbf{l}}}^{\text{gr}}$ is due to the map ${\mathcal{V}}_{{\boldsymbol{n}}}\rightarrow V_{\boldsymbol{x}}\otimes_{{\mathbb{Q}}}{\mathbb{K}}$ of (88) shifting the variables according to $z_{ia}\leadsto z_{ia}x_{ia}$..
4.6. Framed stable critical $K$-theory
Following a suggestion of Shivang Jindal, we restrict the objects (102) and (103) to the stable open subsets $M_{{\boldsymbol{n}},{\boldsymbol{m}}}^{\text{stab}}\subset M_{{\boldsymbol{n}},{\boldsymbol{m}}}$ and $M_{\boldsymbol{x},{\boldsymbol{m}}}^{\text{stab}}\subset M_{\boldsymbol{x},{\boldsymbol{m}}}$, thus obtaining the framed stable critical $K$-theory groups  [^7] Stable framed critical $K$-theory was studied in [47, 48], see also [9, 45] in cohomology.
(108) $K^{{\mathbb{C}}^{*}}(Q,W)^{\text{stab}}_{{\boldsymbol{n}},{\boldsymbol{m}}}\xrightarrow{\iota^{\text{stab}}_{{\boldsymbol{n}},{\boldsymbol{m}}}}K^{{\mathbb{C}}^{*}}(M_{{\boldsymbol{n}},{\boldsymbol{m}}}^{\text{stab}})\cong{\mathcal{V}}_{{\boldsymbol{n}}}\Big/\Theta\left(\frac{1}{{\boldsymbol{\psi}}^{\text{den}}}\right)_{{\boldsymbol{n}}}$
(109) $K(Q^{\text{gr}},W^{\text{gr}})^{\text{stab}}_{\boldsymbol{x},{\boldsymbol{m}}}\xrightarrow{\iota^{\text{gr},\text{stab}}_{\boldsymbol{x},{\boldsymbol{m}}}}K(M_{\boldsymbol{x},{\boldsymbol{m}}}^{\text{stab}})\cong V_{\boldsymbol{x}}\Big/\Theta_{\boldsymbol{x}}^{{\boldsymbol{m}}}$
[…]
The top squares commute, while the bottom ones do not.
Proof.
of Theorem 1.8: We showed in Proposition 4.2 that $\text{Im }\iota_{{\boldsymbol{n}}}={\mathcal{S}}_{{\boldsymbol{n}}}$. Thus, the image of the composition of the top two horizontal maps is the vector space $\Delta_{{\mathbf{l}}}\cdot{\mathcal{S}}_{{\boldsymbol{n}}}$. Mapping further down the two vertical maps, we obtain a vector space isomorphic to
[…]
As in (16), multiplication by $\Delta_{{\mathbf{l}}}$ gives an isomorphism between the LHS above and
${\mathcal{S}}_{{\boldsymbol{n}}}\Big/\left({\mathcal{S}}_{{\boldsymbol{n}}}\cap\Theta\left(\frac{{\boldsymbol{\psi}}^{\text{num}}}{{\boldsymbol{\psi}}^{\text{den}}}\right)_{\boldsymbol{x}}\right)$
precisely as we needed to show. ∎
4.7. Coda
We have just shown that we can obtain a vector space naturally isomorphic to (23) as the image of the composition $\rightarrow\rightarrow\downarrow\downarrow$ in diagram (110).
Lemma 4.8.
The compositions $\rightarrow\rightarrow\downarrow\downarrow$ and $\downarrow\downarrow\rightarrow\rightarrow$ in (110) have the same image.
Proof.
[…]
$\Big[\text{middle }K^{{\mathbb{C}}^{*}}(Q,W)^{\text{stab}}_{{\boldsymbol{n}}}\Big]\rightarrow V_{\boldsymbol{x}}/\Theta_{\boldsymbol{x}}^{{\boldsymbol{m}}}\otimes_{{\mathbb{Q}}}{\mathbb{K}}$
[…]
$\Delta_{{\mathbf{l}}}\Big|_{{\mathbb{C}}^{*}\text{ fixed locus}}=\Delta_{{\mathbf{l}}}^{\text{gr}}\cdot u$
where $u$ is invertible in the ring $V_{\boldsymbol{x}}/\Theta_{\boldsymbol{x}}^{{\boldsymbol{m}}}\otimes_{{\mathbb{Q}}}{\mathbb{K}}$, all compositions
$\Big[\text{leftmost }K^{{\mathbb{C}}^{*}}(Q,W)^{\text{stab}}_{{\boldsymbol{n}}}\Big]\rightarrow V_{\boldsymbol{x}}/\Theta_{\boldsymbol{x}}^{{\boldsymbol{m}}}\otimes_{{\mathbb{Q}}}{\mathbb{K}}$
in diagram (110) have the same image.
Lemma 4.8 implies that we may realize the vector space (23) as the image of the composition $\downarrow\downarrow\rightarrow\rightarrow$ in (110), and we will argue that this composition is more natural than $\rightarrow\rightarrow\downarrow\downarrow$ from a geometric point of view. Indeed, it was proved in [5] that
(111) $\pi\text{ is surjective and }\iota_{\boldsymbol{x},{\boldsymbol{m}}}^{\text{gr},\text{stab}}\text{ is injective}$
using their theory of critical stable envelopes. Therefore, the image of the composition $\downarrow\downarrow\rightarrow\rightarrow$ in (110) is isomorphic to the image of the map $\cdot\Delta_{{\mathbf{l}}}^{\text{gr}}$ therein. Thus
(112) ${\mathcal{S}}_{{\boldsymbol{n}}}\Big/\left({\mathcal{S}}_{{\boldsymbol{n}}}\cap\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}}\right)\cong\left[\Delta^{\text{gr}}_{{\mathbf{l}}}\cdot K(Q^{\text{gr}},W^{\text{gr}})^{\text{stab}}_{{\boldsymbol{m}},\boldsymbol{x}}\right]\otimes_{{\mathbb{Q}}}{\mathbb{K}}$
which is the content of Corollary 1.10. Since the left-hand side of the above equation categorifies the coefficients of the $q$-character, (112) is precisely the categorification and generalization to all $\boldsymbol{\psi}$} of Theorem 1.4: since critical $K$-theory of quivers with potential is a natural replacement of the $K$-theory of the critical locus, then
$K(Q^{\text{gr}},W^{\text{gr}})^{\text{stab}}_{\boldsymbol{x},{\boldsymbol{m}}}\quad\text{is a replacement of}\quad K(M_{\boldsymbol{x},{\boldsymbol{m}}}^{\text{stab}}\cap\text{Crit}(\text{Tr}(W^{\text{gr}})))$
Meanwhile, multiplication by $\Delta_{{\mathbf{l}}}^{\text{gr}}$ is a derived version of imposing the vanishing of $l_{i,s}$ generic linear combinations of paths ending at any vertex $(i,s)$. Therefore, the RHS of (112) is a natural replacement of $K(N_{\boldsymbol{x},{\boldsymbol{m}},{\mathbf{l}}}^{\text{stab}})\otimes_{{\mathbb{Q}}}{\mathbb{K}}$.